### v1

# X-GAN: A Generative AI-Powered Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma

Abstract
Structural changes in main retinal blood vessels serve as critical biomarkers for the onset and progression of glaucoma. Identifying these vessels is vital for vascular modeling yet highly challenging. This paper proposes X-GAN, a generative AI-powered unsupervised segmentation model designed for extracting main blood vessels from Optical Coherence Tomography Angiography (OCTA) images. The process begins with the Space Colonization Algorithm (SCA) to rapidly generate a skeleton of vessels, featuring their radii. By synergistically integrating generative adversarial networks (GANs) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels. Based on this reconstruction, X-GAN achieves nearly 100% segmentation accuracy without relying on labeled data or high-performance computing resources. Also, to address the Issue, data scarity, we introduce GSS-RetVein, a high-definition mixed 2D and 3D glaucoma retinal dataset. GSS-RetVein provides a rigorous benchmark due to its exceptionally clear capillary structures, introducing controlled noise for testing model robustness. Its 2D images feature sharp capillary boundaries, while its 3D component enhances vascular reconstruction and blood flow prediction, supporting glaucoma progression simulations. Experimental results confirm GSS-RetVein’s superiority in evaluating main vessel segmentation compared to existing datasets. Code and dataset are here: https://github.com/VikiXie/SatMar8.
1 Introduction
Glaucoma is one of the primary causes of irreversible blindness and vision impairment [35, 37, 58]. Early detection is crucial, as the disease often progresses silently, with noticeable symptoms appearing only after significant damage has already occurred [21, 22]. Recent studies have identified changes in choroidal blood vessels, especially microvasculature dropout, as promising new biomarkers for monitoring glaucoma progression [41, 54]. These changes can be assessed using the advanced imaging technology of Optical Coherence Tomography Angiography (OCTA) techniques [30, 23]. However, accurately evaluating choroidal vessel density using OCTA is a complex and nontrivial task.
Existing medical imaging analytics in glaucoma research, such as extraction and segmentation of major retinal blood vessels, predominantly depend on computer vision techniques [29, 27, 31, 28, 23, 37, 58, 57, 38, 36, 35, 19]. In recent years, supervised learning segmentation models (SLSMs) [49, 19, 44, 42, 31, 27, 28, 16, 39, 29, 62, 6, 64, 9] have become particularly prominent in this field. However, SLSMs may not be well-suited for detecting choroidal vessels in OCTA images. Supervised learning approaches rely heavily on labeled data, but manually annotating choroidal blood vessels in OCTA images is extremely challenging with current devices [30, 40]. OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [30, 40]. However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [26, 21], and complex capillary structures. Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [21, 22].
To fill the gap, this research introduces a novel unsupervised approach to detecting blood vessels in OCTA images.
Glaucoma researchers have revealed that retinal blood vessels follow biostatistical relationships, meaning that vessel radius variations along branching structures [7, 23]. raising the question of whether segmentation can bypass pixel-based mapping and leverage vessel radius as a key marker. To address this, we propose X-GAN, an unsupervised model that integrates GANs [13] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training. To tackle data scarcity, we introduce GSS-RetVein, a benchmark dataset with clear capillary distribution in 2D and 3D, enhancing segmentation, 3D reconstruction, and blood flow prediction.
In summary, the contributions of this paper are three-fold:
  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images.
  • We propose the GSS-RetVein, a unique and pioneering dataset focused exclusively on retinal vascularity. GSS-RetVein consists of 2D and 3D imaging data, and also preserves intricate capillary details without compromising resolution, ensuring sharp blood vessel boundaries and minimal pixel-induced blurring.
  • Experimental results of X-GAN on GSS-RetVein, OCTA-500 [30] and ROSE [40] for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) cureent models, achieving nearly 100% segmentation accuracy. Also, experimental results of different models on GSS-RetVein have shown the superiority of GSS-RetVein.
2 Related Work
2.1 Datasets for Glaucoma
Many datasets of glaucoma focus on computer vision tasks [37, 58, 57, 38, 36, 35, 28, 40, 30, 63, 52, 53]. For inner retinal blood vessels, datasets like LAG [28], OCTA-500 [30] and ROSE [40], provide various-sizeimages (e.g., 3mm, 6mm). However, these images represent only a small portion of the total dataset. Some datasets [63, 52, 53, 37, 58, 57, 38, 36, 35, 28] offer fundus images, and the capillaries are nearly indistinguishable to the naked eye. There is a shortage of high-quality, intuitive data, such as detailed images of inner 3mm×$\times$3mm retinal vessels, and existing datasets often lack sufficient clarity. Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
Figure 1: The architecture of X-GAN. Red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
2.2 Computer Vision for Glaucoma
Even when data is accessible, challenges in labeling and the subjective variability among doctors may introduce biases or hinder optimal model fitting [37, 58, 57, 38, 36, 35, 19]. Generative AI can compensate for data deficiencies and labeling issues, though it often requires human verification [60] and generated data also face challenges in labeling. However, GAN [13, 66, 20, 65, 24, 61] is an unsupervised model which can solve both issues: data scarity and labeling. GAN-based segmentation methods are generally designed in conjunction with a dedicated segmentation network. For the segmentor part $S$, many pre-trained models like CNNs [27, 31, 28], U-Net [49, 64, 19, 44], ViT [10], Mask R-CNN [16] and MedSAM [39] can be used for final segmentation. Most recent studies from CVPR [50, 20, 2, 8], IEEE Transactions on Medical Imaging [23, 61, 11, 24] and other top conferences or journals [59, 3, 65] also follow this way.
2.3 Motivation
For GAN+SLSM [23, 61, 2, 11, 24, 20], or SLSMs [49, 19, 44, 42, 31, 27, 28, 16, 39, 29, 62, 6, 64, 9], pixel-level representations can be subjective, as different doctors may annotate the details of the same object differently [35, 30, 29]. In OCTA imaging, retinal blood vessels exhibit a uniform color change and have a shape closely resembling a path. They can be represented using coordinates along with road width. Space Colonization Algorithm (SCA) [46] just meets this requirement. Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity. The graph consists of the following components, shown in Eq. 1:
$V=\left\{(x_{i},y_{i},z_{i}|r_{i})\right\}_{i=1}^{N}$ (1)
where ($x_{i}$, $y_{i}$, $z_{i}$) are the spatial coordinates of the vessel centerline points, $r_{i}$ represents the local vessel radius and $N$ means the number of nodes. This structured representation is initially generated using the SCA [46, 32, 25], which models vessel growth based on attraction points and bifurcation principles. However, the initial vessel maps exhibit domain discrepancies in contrast and noise characteristics compared to real OCTA images. To mitigate this gap, we employ a tuned GAN based on CycleGAN [66] as a structural refinement module, aligning vessel representations with real OCTA distributions while preserving vascular topology.
We introduce the Depth-First Search (DFS, a nineteenth century algorithm)-based segmentation approach [55], applying DFS directly to structured (coordinates, radius) data before image rendering. Unlike GAN+SLSM [23, 61, 2, 11, 24, 20], which segments generated images mixed with real data, our method intercepts the GAN generation process, ensuring segmentation aligns with biostatistical vessel radii. The DFS threshold, $R_{min}$, is defined as the minimum radius ratio of main vessels, following established biostatistical principles [22, 12, 5, 43, 7].
Also, we introduce GSS-RetVein, a dataset capturing rich details of both main vessels and capillaries. As shown in Fig. 4, its images (2D, Fig. 4(b)) retain clear capillary edges and tips, unlike ROSE [40] (Fig. 4(d)) and OCTA-500 [30] (Fig. 4(c)). The dense capillaries introduce noise, effectively testing model robustness for SLSMs. This detail aids vascular reconstruction, blood flow prediction, and glaucoma progression simulation by enhancing edge and endpoint learning [30, 29, 62]. Inspired by the Harvard glaucoma dataset series [37, 58, 57, 38, 36, 35], we integrate fairness measures to minimize bias and ensure equitable model performance.
Note: given the model’s simplicity (adjusted GAN + DFS) yet crazy performance (nearly 100% segmentation accuracy), we name it ”X” to highlight its exceptional segmentation ability, similar to how Elon Musk rebranded Twitter as X. Thus, we refer to our model as X-GAN.
3 Methodology: X-GAN
The complete framework of X-GAN is illustrated in Fig. 1 and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
3.1 Vessel Structure Refinement Module
Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$. To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$. This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Eq. 2:
$L_{\text{seg}}=\mathbb{E}_{X_{r}}\left[\|S(X_{r}^{\prime})-S(X_{r})\|_{1}\right]$ (2)
where $S(X)$ extracts vessel masks from OCTA images, enforcing structural consistency during transformation. This constraint ensures that CycleGAN adapts contrast while strictly preserving vessel topology. Consequently, our total loss function is formulated as follows:
$L_{\text{GAN}}$ $=\mathbb{E}_{X_{r}}[D(X_{r})]-\mathbb{E}_{X_{s}}[D(G(X_{s}))]$ (3)
$L_{\text{GP}}$ $=\lambda\mathbb{E}_{\hat{X}}\left[\left(\|\nabla_{\hat{X}}D(\hat{X})\|_{2}-1\right)^{2}\right]$ (4)
$L_{\text{total}}$ $=L_{\text{GAN}}+\lambda_{\text{GP}}L_{\text{GP}}+\lambda_{\text{seg}}L_{\text{seg}}$ (5)
where Eq. 3 is the Wasserstein Adversarial Loss, Eq. 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Eq. 5 is the Structural Consistency via Segmentation Loss. Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
3.2 Segmentor
Rather than utilizing pixel-wise segmentation maps, we directly extract primary vessels from the $(x_{i}^{G},y_{i}^{G},z_{i}^{G}|r_{i}^{G})$ representation using DFS based graph traversal, which effectively isolates large vessels while filtering out capillaries.
First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity. A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Eq. 6:
$w(e_{ij})=r_{i}$ (6)
where larger vessels are prioritized in the traversal process. To extract primary vessels, we apply radius thresholding and DFS traversal. For radius filtering, we define a minimum main vessel radius $R_{min}$ and retain only Eq. 7:
$V_{\text{main}}=\left\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\min}*r_{max})\right\}$ (7)
For DFS traversal, it consists of three steps: (1) select the optic disc region as the root node; (2) recursively traverse the largest connected component using DFS; (3) stop when no further vessel segments satisfy the radius constraint. When all three steps finished, construct final extracted vessel structure $G_{main}$. This approach effectively retains only the primary vascular structure, removing small capillaries and noise. This method can be summarized as the following Algorithm 1.
Algorithm 1 : Primary Vessel Extraction via DFS
1:Vessel structure $V=\{(x_{i},y_{i},z_{i}|r_{i})\}_{i=1}^{N}$
2:Extracted primary vessel graph $G_{\text{main}}$
3:Step 1: Vessel Graph Construction
4:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V$ do
5:     Identify neighboring vessel points to form edges $E$
6:     Assign edge weights: $w(e_{ij})=r_{i}$
7:end for
8:Step 2: Vessel Filtering
9:Apply radius threshold $R_{\text{min}}$:
10:$V_{\text{main}}=\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\text{min}}*r_{max})\}$
11:Remove weakly connected components
12:Step 3: DFS for Vessel Traversal
13:Initialize DFS from optic disc region
14:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V_{\text{main}}$ do
15:     Recursively explore vessel network using DFS
16:     Expand along larger vessels
17:end for
18:Step 4: Output Primary Vessel Graph return $G_{\text{main}}$
The key aspect of Algorithm 1 is the threshold $R_{min}$. Instead of the final image, we analyze intermediate outputs $S(X_{r})$, CSV data (Fig. 2(e)), and GAN-derived radius statistics (Fig. 2(d)). For a sample image (Fig. 2(a)), we computed a Mean (85.31) and Standard Deviation (88.67), determining $R_{min}$ as 0.1968 $R_{max}$. Across datasets, values were 0.2011 (GSS-RetVein), 0.2102 (OCTA-500), and 0.1982 (ROSE), standardized to 0.2 for consistency. Ophthalmology studies [22, 12, 5, 43, 7] confirm 18-26% capillary coverage in OCTA images, validating this threshold.
(a) GSS-RetVein
(b) 3D Visualization
(c) Statistical Distribution
(d) Radius Statistics
(e) Vascular Data Format (Part)
Figure 2: 3D visualization of the pixel distribution of one sample, and its Gaussian statistical distribution.
[…]
4 Dataset and Implement
4.1 Dataset
4.1.1 GSS-RetVein
GSS-RetVein is a unique and pioneering dataset focused exclusively on retinal vascularity. The dataset features an exceptionally clear distribution of capillaries. It comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race. We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female. Racial distribution reflects local glaucoma clinic demographics: 80.4% white, 14.6% Black, with Asians and other minority groups comprising the remainder. Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
GSS-RetVein’s original images come from the latest Intalight device [^1] https://intalight.com/, offering superior detail compared to OPTOVUE [^2] https://www.visionix.com/ (Fig. 3). As shown in Fig. 3(a) and Fig. 3(c), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy. In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [62]. For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [29], making Intalight data ideal for both tasks.
(a) Intalight
(b) OPTOVUE
(c) Intalight
(d) OPTOVUE
Figure 3: Comparison of data collected by different devices for the same patient. Among them, Fig. 3(a) and Fig. 3(b) are images of deep angio, and Fig. 3(c) and Fig. 3(d) are images of superficial angio. They all are images of inner retina vessel.
For GSS-RetVein, we eliminate red and green baselines from Intalight device images (Fig. 4(a)) through a multi-step process. First, we apply a dark channel algorithm [15] to remove light artifacts, followed by unsharp masking [51] to enhance high-frequency details. The final images (Fig. 4(b)) undergo manual verification for quality assurance. As shown in Fig. 4, while main vessels remain visible across datasets, magnification reveals that OCTA-500 [30] and ROSE [40] blur capillaries and vessel ends. Even the unprocessed baseline image (Fig. 4(a)) offers superior vessel clarity compared to other datasets (Fig. 4(d), Fig. 4(c)).
(a) Baseline
(b) GSS-RetVein
(c) OCTA-500
(d) ROSE
Figure 4: Comparison of GSS-RetVein images with other datasets.
For the representation of 3D imaging of GSS-RetVein, as shown in Fig. 5, capillaries stand out from the thicker main blood vessels by attaching to them and narrowing at their ends (Fig. 5(c)). The elaborate architecture of blood vessels and the intertwining capillaries contribute to an extremely intricate vascular layout. 3D imaging part of GSS-RetVein is formatted as a JSON file, which includes 3D X, Y, and Z coordinates and the corresponding radius pixel distribution values. For annotation (2D, image format: PNG), our team includes two glaucoma specialists who handle labeling, verification, and manual evaluation, utilizing the Labelme [^3] https://github.com/wkentaro/labelme.
(a) Top View
(b) Front View
(c) Detail
Figure 5: 3D imaging of retinal vessel shown from various angles.
In general, GSS-RetVein is a dataset that combines 2D and 3D images with extremely high image resolution, and is particularly dedicated to retinal vascular segmentation.
Model Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
U-Net [49] 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
U-Net++ [64] 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
Attention U-Net [44] 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
Mask R-CNN [16] 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
YOLOv11-x [18] 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
MedSAM [39] 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
CauSSL [42] 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
UniverSeg [6] 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
S2VNet [9] 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
Tyche [45] 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
X-GAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
X-GAN (MAX) 99.89 99.94 99.66 99.83 99.73 99.86 99.80 99.89 99.57 99.78
Table 1: Comparison X-GAN With other models on GSS-RetVein, OCTA-500 and ROSE. ($\times$100%)
Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki. Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
4.1.2 OCTA-500
OCTA-500 [30] focuses on retinal blood vessel research, providing 6mm×$\times$6mm (300 images, ID range 10001 to 10300) and 3mm×$\times$3mm (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies. It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
4.1.3 ROSE
ROSE [40] is an open-source collection designed for retinal blood vessel segmentation using OCTA images. It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a 3mm×$\times$3mm foveal-centered area with a resolution of 304$\times$304 pixels, providing both centerline-level and pixel-level vessel annotations. ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
4.2 Implement
4.2.1 Hyperparameter
For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70). The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$. X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation. Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
The hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
4.2.2 Evaluation Metric
We use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth: $\text{IoU}=|P\cap G|/|P\cup G|$. This metric ranges from 0 to 1, with 1 indicating perfect segmentation. Additionally, we use the Dice Coefficient, another metric for segmentation accuracy that calculates overlap between the prediction and ground truth: $\text{Dice}=2\times|P\cap G|/(|P|+|G|)$. In both formulas, $P$ represents predicted vessel pixels, $G$ the ground truth vessel pixels, $|P\cap G|$ the count of correctly predicted vessel pixels, and $|P\cup G|$ the combined set of ground truth and predicted vessel pixels. All in all, IoU is used to evaluate segmentation accuracy of the main vessels, while the Dice Coefficient assesses accuracy at finer vessel ends relative to the main vessels. For vascular structural enhancement, we use SSIM (Structural Similarity Index) and MSE (Mean Squared Error). SSIM (range: -1 to 1) assesses similarity based on luminance, contrast, and structure, while MSE quantifies pixel-wise differences, with lower values indicating higher similarity. While SSIM aligns with human perception, PSNR and MSE focus on absolute pixel differences.
(a) Comparison of models, trained on GSS-RetVein and OCTA and tested on ROSE.
(b) Comparison of models, trained on ROSE and tested on GSS-RetVein and OCTA.
(c) Comparison of models, trained on GSS-RetVein or OCTA and tested on each other.
Figure 6: Conducting mixed testing on different models using mixed datasets. ($\times$100%)
5 Experimental Result
Model Dataset
GSS-RetVein OCTA-500 ROSE
6mm 3mm 6mm 3mm 3mm
SSIM MSE SSIM MSE SSIM MSE SSIM MSE SSIM MSE
CycleGAN [66] 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
CycleGAN (ours) 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
Table 2: Comparison vessel structure refinement images with real OCTA images. ($\times$100%)
5.1 Comparative Experiment
As presented in Tab. 1, experimental findings demonstrate that while models such as MedSAM, U-Net++, and YOLOv11-x exhibit competitive performance, X-GAN consistently outperforms all competing models across varying dataset resolutions, achieving the highest scores and excelling in retinal vessel segmentation, where both of the evaluation matrixs are even close to 100%, highlighting its superior segmentation accuracy and underscoring its robustness in fine-grained vascular structures segmentation.
As shown in Fig. 6(a), given the comparable sizes of GSS-RetVein and OCTA-500 (retinal vessel subset) and the smaller size of ROSE, we trained models on the first two and tested on the latter. X-GAN consistently achieved the highest IoU and Dice scores on ROSE, outperforming S2VNet, CauSSL, and MedSAM. To further assess generalization, we trained models on ROSE and tested them on GSS-RetVein and OCTA-500 (Fig. 6(b)), where X-GAN again achieved the highest scores, followed by MedSAM and CauSSL. Finally, Fig. 6(c) evaluates models across different scanning resolutions, with X-GAN consistently leading in segmentation accuracy. While MedSAM and UniverSeg performed well, U-Net and its variants showed strong but relatively lower performance. These results highlight GSS-RetVein’s high quality and X-GAN’s superior segmentation capability.
Note: the highest indicator is shown in bold, while the second highest is marked with an underline. We do not consider X-GAN (MAX) in all Tabs. 1, 6 and 3.
5.2 Ablation Experiment
For baseline GAN model, GAN [13] and CycleGAN [66] are selected. For baseline segmentor, U-Net [49], U-Net++ [64], Attention U-Net [44] and MedSAM [39] are selected. For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy. As shown in Tab. 3, since main blood vessels make up a small proportion, the value calculated is also relatively low. When the minimum radius ratio is found, like X-GAN (MAx), aligning with biostatistical properties [22, 12, 5, 43, 7]—segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
Model Dataset
GSS-RetVein
6mm 3mm
IoU Dice IoU Dice
GAN [13] + U-Net [49] 92.89 96.31 92.44 96.07
GAN [13] + U-Net++ [64] 93.62 96.70 92.87 96.30
GAN [13] + Attention U-Net [44] 93.56 96.67 93.12 96.44
GAN [13] + MedSAM [39] 93.73 96.76 92.76 96.24
GAN [13] + DFS ($R_{min}$=0.2) 95.81 97.86 95.24 97.56
CycleGAN [66] + U-Net [49] 93.44 96.61 93.42 96.60
CycleGAN [66] + U-Net++ [64] 93.72 96.76 93.62 96.70
CycleGAN [66] + Attention U-Net [44] 93.79 96.80 93.45 96.61
CycleGAN [66] + MedSAM [39] 94.02 96.92 94.37 97.10
CycleGAN [66] + DFS ($R_{min}$=0.2) 96.23 98.08 96.42 98.18
CycleGAN (ours) + U-Net [49] 94.56 97.20 94.52 97.18
CycleGAN (ours) + U-Net++ [64] 94.98 97.43 94.36 97.10
CycleGAN (ours) + Attention U-Net [44] 95.06 97.47 95.12 97.50
CycleGAN (ours) + MedSAM [39] 95.33 97.61 95.21 97.55
X-GAN 99.41 99.71 98.66 99.33
X-GAN (MAX) 99.89 99.94 99.66 99.83
Table 3: Ablation experiment of X-GAN. ($\times$100%)
As shown in Fig. 7(a), the optimal choice of $R_{min}$ (0.2) enables X-GAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%. When $R_{min}=1$, X-GAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold. When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
(a) $R_{min}$
(b) The Number of Nodes $N$
Figure 7: Ablation experiments of key parameters. ($\times$100%)
Also, as shown in Eq. 1 and Fig. 7(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy. Given that the R values of X-GAN and X-GAN (MAX) are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar. Consequently, their curves largely overlap, with X-GAN (MAX) exhibiting a slight increase.
5.3 Evaluation of Vessel Structure Refinement
The comparison between CycleGAN [66] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%. CycleGAN (ours) consistently outperforms CycleGAN, achieving higher SSIM and lower MSE across all datasets. The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
Human Evaluation: the medical team verified the corrected images. Luckily, since this study only aimed to segment only the main blood vessels, the small size of the capillaries caused errors in radius generation, which affected the accuracy of the indicators.
Indicator SLSMs DFS (ours)
GPU $\checkmark$ $\times$
Annotation $\checkmark$ $\times$
Training $\checkmark$ $\times$
Explainability $\times$ $\checkmark$
2D and 3D $\times$ $\checkmark$
Generalization $\times$ -
Table 4: Comparison of performance indicators between DFS (our) and SLSMs. ($\times$100%) (- indicates strong performance for a specific class.)
(a) Segmentation of Main Retinal Vessels (2D)
(b) Original Image (2D)
(c) Original Image (3D)
(d) Segmentation (2D)
(e) Segmentation (3D)
Figure 8: Segmentation results of X-GAN on GSS-RetVein (2D and 3D).
5.4 Efficiency Evaluation
All in all, as shown in Tab. 4, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity. It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
What is more, in more detail, the advantages of our approach are as follows in Tab. 5, our segmentor has demonstrated unsurpassed superiority.
Model Details
Inference Time Parameters FLOPs GPU Utilization
U-Net [49] 21ms 31M 98 low
U-Net++ [64] 62ms 76M 304 high
Attention U-Net [44] 83ms 62M 215 slightly high
MedSAM [39] 425ms 109M 1250 extremely high
DFS (ours) 0.018ms 0M 0.00028 -
DFS (ours, GPU based) $\leq$ 0.01ms 0M 0.00028 extremely low
Table 5: Comparison of performance details between DFS (our) and SLSMs.
5.5 Visualization
As shown in Fig. 8, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis. In Fig. 8(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow. DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries. Extending this to 3D, we achieve precise main vessel extraction, with Fig. 8(e) preserving vessel integrity and offering a clearer structure than the original Fig. 8(c).
5.6 Analysis
Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Tabs. 1, 6 and 3). Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [22, 12, 5, 43, 7]. We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results. Studies [34, 56] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models. There are many examples: Faster R-CNN [48] and Mask R-CNN [16], YOLOv3 [47], YOLOv4 [4] and YOLOv5 [17], DeepSeek [14, 33] and ChatGPT [1].
To the best of our knowledge, X-GAN is the best model currently, tailored to segment OCTA retinal main vessel.
6 Conclusion
In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs. We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation. Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability. Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
Acknowledgements
This research is sponsored by NASA 80NSSC22K0144, UTSW GMO No. 241213, and NIH 1R01AG083179-01.

### v3

@@
-X-GAN: A Generative AI-Powered Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma
+X-GAN: A Generative AI-Powered Unsupervised Model for Main Vessel Segmentation of Glaucoma Screening
 Abstract
@@
 The process begins with the Space Colonization Algorithm (SCA) to rapidly generate a skeleton of vessels, featuring their radii.
-By synergistically integrating generative adversarial networks (GANs) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels.
+By synergistically integrating the generative adversarial network (GAN) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels.
 Based on this reconstruction, X-GAN achieves nearly 100% segmentation accuracy without relying on labeled data or high-performance computing resources.
-Also, to address the Issue, data scarity, we introduce GSS-RetVein, a high-definition mixed 2D and 3D glaucoma retinal dataset.
-GSS-RetVein provides a rigorous benchmark due to its exceptionally clear capillary structures, introducing controlled noise for testing model robustness.
-Its 2D images feature sharp capillary boundaries, while its 3D component enhances vascular reconstruction and blood flow prediction, supporting glaucoma progression simulations.
-Experimental results confirm GSS-RetVein’s superiority in evaluating main vessel segmentation compared to existing datasets.
-Code and dataset are here: https://github.com/VikiXie/SatMar8.
+Experimental results confirm X-GAN’s superiority in evaluating main vessel segmentation compared to existing deep learning models.
+Code is here: https://github.com/VikiXie/SatMar8.
+Keywords:
+Medical Imaging Unsupervised Learning Glaucoma Pathological Segmentation Generative AI
 1 Introduction
-Glaucoma is one of the primary causes of irreversible blindness and vision impairment [35, 37, 58].
-Early detection is crucial, as the disease often progresses silently, with noticeable symptoms appearing only after significant damage has already occurred [21, 22].
-Recent studies have identified changes in choroidal blood vessels, especially microvasculature dropout, as promising new biomarkers for monitoring glaucoma progression [41, 54].
-These changes can be assessed using the advanced imaging technology of Optical Coherence Tomography Angiography (OCTA) techniques [30, 23].
-However, accurately evaluating choroidal vessel density using OCTA is a complex and nontrivial task.
-Existing medical imaging analytics in glaucoma research, such as extraction and segmentation of major retinal blood vessels, predominantly depend on computer vision techniques [29, 27, 31, 28, 23, 37, 58, 57, 38, 36, 35, 19].
+Glaucoma is a leading cause of irreversible blindness, often progressing silently until significant vision loss occurs [40, 37, 38, 42, 30].
+Recent research highlights choroidal microvasculature dropout as a potential biomarker for disease progression [42].
+Optical Coherence Tomography Angiography (OCTA) enables visualization of these vascular changes, but accurately assessing choroidal vessel density remains a challenging task [1, 17].
+Existing medical imaging analytics in glaucoma research predominantly depend on computer vision techniques [2, 45, 44, 43, 17, 37, 38, 36, 39, 41, 40, 46].
 In recent years, supervised learning segmentation models (SLSMs) [10, 46, 12, 13, 44, 45, 43, 9, 8, 2, 3, 14, 7, 15] have become particularly prominent in this field.
 However, SLSMs may not be well-suited for detecting choroidal vessels in OCTA images.
-Supervised learning approaches rely heavily on labeled data, but manually annotating choroidal blood vessels in OCTA images is extremely challenging with current devices [30, 40].
+They rely heavily on labeled data, but manually annotating choroidal blood vessels in OCTA images is extremely challenging with current devices [1, 4].
 OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [1, 4].
-However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [26, 21], and complex capillary structures.
+However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [42], and complex capillary structures.
 Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [42, 30].
-To fill the gap, this research introduces a novel unsupervised approach to detecting blood vessels in OCTA images.
 Glaucoma researchers have revealed that retinal blood vessels follow biostatistical relationships, meaning that vessel radius variations along branching structures [32, 17].
@@
 To address this, we propose X-GAN, an unsupervised model that integrates GANs [34] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training.
-To tackle data scarcity, we introduce GSS-RetVein, a benchmark dataset with clear capillary distribution in 2D and 3D, enhancing segmentation, 3D reconstruction, and blood flow prediction.
 In summary, the contributions of this paper are three-fold:
-  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images.
-  • We propose the GSS-RetVein, a unique and pioneering dataset focused exclusively on retinal vascularity.
-GSS-RetVein consists of 2D and 3D imaging data, and also preserves intricate capillary details without compromising resolution, ensuring sharp blood vessel boundaries and minimal pixel-induced blurring.
-  • Experimental results of X-GAN on GSS-RetVein, OCTA-500 [30] and ROSE [40] for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) cureent models, achieving nearly 100% segmentation accuracy.
-Also, experimental results of different models on GSS-RetVein have shown the superiority of GSS-RetVein.
+  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images without annotations.
+  • We eliminate the reliance on synthetic label supervision mechanisms (SLSMs) by introducing a novel segmentation strategy based on vessel radius thresholds and depth-first search (DFS), enabling accurate main vessel extraction without any pixel-level annotations.
+  • Experimental results of X-GAN for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) current models, achieving nearly 100% segmentation accuracy.
 2 Related Work
-2.1 Datasets for Glaucoma
-Many datasets of glaucoma focus on computer vision tasks [37, 58, 57, 38, 36, 35, 28, 40, 30, 63, 52, 53].
+2.1 Medical Imaging in Glaucoma
+Many datasets of glaucoma focus on computer vision tasks [37, 38, 36, 39, 41, 40, 43, 4, 1, 53, 54, 52], especially in medical image processing.
 For inner retinal blood vessels, datasets like LAG [43], OCTA-500 [1] and ROSE [4], provide various-sizeimages (e.g., 3mm, 6mm).
@@
 Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
-Figure 1: The architecture of X-GAN. Red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
-2.2 Computer Vision for Glaucoma
-Even when data is accessible, challenges in labeling and the subjective variability among doctors may introduce biases or hinder optimal model fitting [37, 58, 57, 38, 36, 35, 19].
-Generative AI can compensate for data deficiencies and labeling issues, though it often requires human verification [60] and generated data also face challenges in labeling.
-However, GAN [13, 66, 20, 65, 24, 61] is an unsupervised model which can solve both issues: data scarity and labeling.
-GAN-based segmentation methods are generally designed in conjunction with a dedicated segmentation network.
-For the segmentor part $S$, many pre-trained models like CNNs [27, 31, 28], U-Net [49, 64, 19, 44], ViT [10], Mask R-CNN [16] and MedSAM [39] can be used for final segmentation.
-Most recent studies from CVPR [50, 20, 2, 8], IEEE Transactions on Medical Imaging [23, 61, 11, 24] and other top conferences or journals [59, 3, 65] also follow this way.
+2.2 AI for Glaucoma
+Even with accessible data, labeling remains challenging due to annotation costs and inter-observer variability, introducing bias and hindering model performance [37, 38, 36, 39, 41, 40, 46].
+Generative AI can alleviate data scarcity and labeling constraints, though human oversight is often required [50].
+GANs [34, 35, 18, 24, 27, 25], as unsupervised models, address both issues by generating realistic samples without manual labels.
+In this framework, segmentation is typically achieved by coupling the GAN with a dedicated segmentor $S$, which can be implemented using pre-trained CNNs [45, 44, 43], U-Net variants [10, 7, 46, 12], ViT [22], Mask R-CNN [9], or MedSAM [8], enabling end-to-end vessel synthesis and segmentation.
 2.3 Motivation
-For GAN+SLSM [23, 61, 2, 11, 24, 20], or SLSMs [49, 19, 44, 42, 31, 27, 28, 16, 39, 29, 62, 6, 64, 9], pixel-level representations can be subjective, as different doctors may annotate the details of the same object differently [35, 30, 29].
-In OCTA imaging, retinal blood vessels exhibit a uniform color change and have a shape closely resembling a path.
-They can be represented using coordinates along with road width.
-Space Colonization Algorithm (SCA) [46] just meets this requirement.
+In traditional models such as GAN+SLSM [17, 25, 20, 26, 27, 18] and standalone SLSMs [10, 46, 12, 13, 44, 45, 43, 9, 8, 2, 3, 14, 7, 15], pixel-level segmentations are prone to inter-observer variability [40, 1, 2].
+However, in OCTA, retinal vessels have consistent visual traits, uniform grayscale gradients and path-like shapes, making them suitable for coordinate-based modeling with associated widths.
+We adopt the Space Colonization Algorithm (SCA) [49] to represent vessels as structured graphs, enhancing topological coherence and connectivity.
+To bridge the domain gap between initial vessel maps and real OCTA data, we apply a CycleGAN-based refinement module [35], aligning generated structures with realistic image characteristics while preserving topology.
+Unlike prior methods that segment post-rendered images, we introduce a DFS-based segmentation approach [51] applied directly to graph-structured (coordinates, radius) data.
+This intercepts the generation process to ensure anatomical fidelity, governed by a radius threshold $R_{min}$ derived from biostatistical vessel metrics [30, 31, 33, 32].
+3 Methodology
+Figure 1: The architecture of X-GAN where red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
+The complete framework of X-GAN is illustrated in Fig. 1 and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
+3.1 Vessel Structure Refinement Module
 Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity.
-The graph consists of the following components, shown in Eq.
-1:
+The graph consists of the following components, shown in Equation 1:
 $V=\left\{(x_{i},y_{i},z_{i}|r_{i})\right\}_{i=1}^{N}$ (1)
@@
 To mitigate this gap, we employ a tuned GAN based on CycleGAN [35] as a structural refinement module, aligning vessel representations with real OCTA distributions while preserving vascular topology.
-We introduce the Depth-First Search (DFS, a nineteenth century algorithm)-based segmentation approach [55], applying DFS directly to structured (coordinates, radius) data before image rendering.
-Unlike GAN+SLSM [23, 61, 2, 11, 24, 20], which segments generated images mixed with real data, our method intercepts the GAN generation process, ensuring segmentation aligns with biostatistical vessel radii.
-The DFS threshold, $R_{min}$, is defined as the minimum radius ratio of main vessels, following established biostatistical principles [22, 12, 5, 43, 7].
-Also, we introduce GSS-RetVein, a dataset capturing rich details of both main vessels and capillaries.
-As shown in Fig.
-4, its images (2D, Fig.
-4(b)) retain clear capillary edges and tips, unlike ROSE [40] (Fig.
-4(d)) and OCTA-500 [30] (Fig.
-4(c)).
-The dense capillaries introduce noise, effectively testing model robustness for SLSMs.
-This detail aids vascular reconstruction, blood flow prediction, and glaucoma progression simulation by enhancing edge and endpoint learning [30, 29, 62].
-Inspired by the Harvard glaucoma dataset series [37, 58, 57, 38, 36, 35], we integrate fairness measures to minimize bias and ensure equitable model performance.
-Note: given the model’s simplicity (adjusted GAN + DFS) yet crazy performance (nearly 100% segmentation accuracy), we name it ”X” to highlight its exceptional segmentation ability, similar to how Elon Musk rebranded Twitter as X.
-Thus, we refer to our model as X-GAN.
-3 Methodology: X-GAN
-The complete framework of X-GAN is illustrated in Fig. 1 and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
-3.1 Vessel Structure Refinement Module
 Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$.
 To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$.
-This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Eq.
-2:
+This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Equation 2:
 $L_{\text{seg}}=\mathbb{E}_{X_{r}}\left[\|S(X_{r}^{\prime})-S(X_{r})\|_{1}\right]$ (2)
@@
 $L_{\text{total}}$ $=L_{\text{GAN}}+\lambda_{\text{GP}}L_{\text{GP}}+\lambda_{\text{seg}}L_{\text{seg}}$ (5)
-where Eq.
-3 is the Wasserstein Adversarial Loss, Eq.
-4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Eq.
-5 is the Structural Consistency via Segmentation Loss.
+where Equation 3 is the Wasserstein Adversarial Loss, Equation 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Equation 5 is the Structural Consistency via Segmentation Loss.
 Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
@@
 First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity.
-A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Eq.
-6:
+A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Equation 6:
 $w(e_{ij})=r_{i}$ (6)
@@
 To extract primary vessels, we apply radius thresholding and DFS traversal.
-For radius filtering, we define a minimum main vessel radius $R_{min}$ and retain only Eq.
-7:
+For radius filtering, we define the ratio of the smallest main vessel radius to the largest main vessel radius $R_{min}$ and retain only Equation 7:
 $V_{\text{main}}=\left\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\min}*r_{max})\right\}$ (7)
@@
 This approach effectively retains only the primary vascular structure, removing small capillaries and noise.
-This method can be summarized as the following Algorithm 1.
-Algorithm 1 : Primary Vessel Extraction via DFS
-1:Vessel structure $V=\{(x_{i},y_{i},z_{i}|r_{i})\}_{i=1}^{N}$
-2:Extracted primary vessel graph $G_{\text{main}}$
-3:Step 1: Vessel Graph Construction
-4:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V$ do
-5:     Identify neighboring vessel points to form edges $E$
-6:     Assign edge weights: $w(e_{ij})=r_{i}$
-7:end for
-8:Step 2: Vessel Filtering
-9:Apply radius threshold $R_{\text{min}}$:
-10:$V_{\text{main}}=\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\text{min}}*r_{max})\}$
-11:Remove weakly connected components
-12:Step 3: DFS for Vessel Traversal
-13:Initialize DFS from optic disc region
-14:for all vessel point $(x_{i},y_{i},z_{i}|r_{i})\in V_{\text{main}}$ do
-15:     Recursively explore vessel network using DFS
-16:     Expand along larger vessels
-17:end for
-18:Step 4: Output Primary Vessel Graph return $G_{\text{main}}$
-The key aspect of Algorithm 1 is the threshold $R_{min}$.
-Instead of the final image, we analyze intermediate outputs $S(X_{r})$, CSV data (Fig.
-2(e)), and GAN-derived radius statistics (Fig.
-2(d)).
-For a sample image (Fig.
-2(a)), we computed a Mean (85.31) and Standard Deviation (88.67), determining $R_{min}$ as 0.1968 $R_{max}$.
-Across datasets, values were 0.2011 (GSS-RetVein), 0.2102 (OCTA-500), and 0.1982 (ROSE), standardized to 0.2 for consistency.
-Ophthalmology studies [22, 12, 5, 43, 7] confirm 18-26% capillary coverage in OCTA images, validating this threshold.
-(a) GSS-RetVein
-(b) 3D Visualization
-(c) Statistical Distribution
-(d) Radius Statistics
-(e) Vascular Data Format (Part)
-Figure 2: 3D visualization of the pixel distribution of one sample, and its Gaussian statistical distribution.
-[…]
-4 Dataset and Implement
+4 Dataset and Implementation
 4.1 Dataset
-4.1.1 GSS-RetVein
-GSS-RetVein is a unique and pioneering dataset focused exclusively on retinal vascularity.
-The dataset features an exceptionally clear distribution of capillaries.
-It comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race.
+4.1.1 OCTA-500
+OCTA-500 [1] focuses on retinal blood vessel research, providing 6mm×$\times$6mm (300 images, ID range 10001 to 10300) and 3mm×$\times$3mm (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies.
+It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
+4.1.2 ROSE
+ROSE [4] is an open-source collection designed for retinal blood vessel segmentation using OCTA images.
+It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a 3mm×$\times$3mm foveal-centered area with a resolution of 304$\times$304 pixels, providing both centerline-level and pixel-level vessel annotations.
+ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
+(a) UTSW Data
+(b) OCTA-500
+(c) ROSE
+Figure 2: Comparison of UTSW with Other Datasets
+4.1.3 UTSW
+our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki.
+Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
+As show in Fig.
+2, original images come from the latest Intalight device [^1] https://intalight.com/, which is currently the most advanced equipment.
+As shown in Fig.
+2(a), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy.
+UTSW comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race.
 We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female.
@@
 Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
-GSS-RetVein’s original images come from the latest Intalight device [^1] https://intalight.com/, offering superior detail compared to OPTOVUE [^2] https://www.visionix.com/ (Fig.
-3).
-As shown in Fig.
-3(a) and Fig.
-3(c), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy.
 In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [3].
 For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [2], making Intalight data ideal for both tasks.
-(a) Intalight
-(b) OPTOVUE
-(c) Intalight
-(d) OPTOVUE
-Figure 3: Comparison of data collected by different devices for the same patient.
-Among them, Fig.
-3(a) and Fig.
-3(b) are images of deep angio, and Fig.
-3(c) and Fig.
-3(d) are images of superficial angio.
-They all are images of inner retina vessel.
-For GSS-RetVein, we eliminate red and green baselines from Intalight device images (Fig.
-4(a)) through a multi-step process.
-First, we apply a dark channel algorithm [15] to remove light artifacts, followed by unsharp masking [51] to enhance high-frequency details.
-The final images (Fig.
-4(b)) undergo manual verification for quality assurance.
-As shown in Fig.
-4, while main vessels remain visible across datasets, magnification reveals that OCTA-500 [30] and ROSE [40] blur capillaries and vessel ends.
-Even the unprocessed baseline image (Fig.
-4(a)) offers superior vessel clarity compared to other datasets (Fig.
-4(d), Fig.
-4(c)).
-(a) Baseline
-(b) GSS-RetVein
-(c) OCTA-500
-(d) ROSE
-Figure 4: Comparison of GSS-RetVein images with other datasets.
-For the representation of 3D imaging of GSS-RetVein, as shown in Fig.
-5, capillaries stand out from the thicker main blood vessels by attaching to them and narrowing at their ends (Fig.
-5(c)).
-The elaborate architecture of blood vessels and the intertwining capillaries contribute to an extremely intricate vascular layout.
-3D imaging part of GSS-RetVein is formatted as a JSON file, which includes 3D X, Y, and Z coordinates and the corresponding radius pixel distribution values.
-For annotation (2D, image format: PNG), our team includes two glaucoma specialists who handle labeling, verification, and manual evaluation, utilizing the Labelme [^3] https://github.com/wkentaro/labelme.
-(a) Top View
-(b) Front View
-(c) Detail
-Figure 5: 3D imaging of retinal vessel shown from various angles.
-In general, GSS-RetVein is a dataset that combines 2D and 3D images with extremely high image resolution, and is particularly dedicated to retinal vascular segmentation.
+Our team includes two attending glaucoma physicians who are responsible for the detailed labeling of the main retinal blood vessels using the tool Labelme [^2] https://github.com/wkentaro/labelme.
+4.2 Implementation
+4.2.1 Hyperparameter
+For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70).
+The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$.
+X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation.
+Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$) and flipping.
+4.2.2 Evaluation Matrix
+We use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth.
+This metric ranges from 0 to 1, with 1 indicating perfect segmentation.
+Additionally, we use the Dice Coefficient, another metric for segmentation accuracy that calculates overlap between the prediction and ground truth.
+All in all, IoU is used to evaluate segmentation accuracy of the main vessels, while the Dice Coefficient assesses accuracy at finer vessel ends relative to the main vessels.
+For vascular structural enhancement, we use Structural Similarity Index (SSIM) and Mean Squared Error (MSE).
+SSIM (range: -1 to 1) assesses similarity based on luminance, contrast, and structure, while MSE quantifies pixel-wise differences, with lower values indicating higher similarity.
+While SSIM aligns with human perception, MSE focus on absolute pixel differences.
+4.2.3 Experimental Setup
+We selected the following deep learning models as reference benchmarks: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [13], UniverSeg [14], S2VNet [15] and Tyche [16].
+These models are leaders in the fields of image segmentation and medical image processing.
+For their parameter adjustment, we use Optuna [^3] https://optuna.org/ to optimize their learning green, decay and other hyperparameters.
+For the training and testing of the model, we use ten cross-validation methods.
+The evaluation indicators are finally averaged.
+4.2.4 Hardware
+The hardware for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
+5 Experimental Result
+5.1 Comparative Experiment
+Table 1: Comparison X-GAN with Other Models on Different Datasets ($\times$100%)
 Model Dataset
-GSS-RetVein OCTA-500 ROSE
+UTSW OCTA-500 ROSE
 6mm 3mm 6mm 3mm 3mm
@@
 X-GAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
-X-GAN (MAX) 99.89 99.94 99.66 99.83 99.73 99.86 99.80 99.89 99.57 99.78
-Table 1: Comparison X-GAN With other models on GSS-RetVein, OCTA-500 and ROSE. ($\times$100%)
-Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki.
-Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
-4.1.2 OCTA-500
-OCTA-500 [30] focuses on retinal blood vessel research, providing 6mm×$\times$6mm (300 images, ID range 10001 to 10300) and 3mm×$\times$3mm (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies.
-It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
-4.1.3 ROSE
-ROSE [40] is an open-source collection designed for retinal blood vessel segmentation using OCTA images.
-It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a 3mm×$\times$3mm foveal-centered area with a resolution of 304$\times$304 pixels, providing both centerline-level and pixel-level vessel annotations.
-ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
-4.2 Implement
-4.2.1 Hyperparameter
-For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70).
-The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$.
-X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation.
-Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
-The hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
-4.2.2 Evaluation Metric
-We use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth: $\text{IoU}=|P\cap G|/|P\cup G|$.
-This metric ranges from 0 to 1, with 1 indicating perfect segmentation.
-Additionally, we use the Dice Coefficient, another metric for segmentation accuracy that calculates overlap between the prediction and ground truth: $\text{Dice}=2\times|P\cap G|/(|P|+|G|)$.
-In both formulas, $P$ represents predicted vessel pixels, $G$ the ground truth vessel pixels, $|P\cap G|$ the count of correctly predicted vessel pixels, and $|P\cup G|$ the combined set of ground truth and predicted vessel pixels.
-All in all, IoU is used to evaluate segmentation accuracy of the main vessels, while the Dice Coefficient assesses accuracy at finer vessel ends relative to the main vessels.
-For vascular structural enhancement, we use SSIM (Structural Similarity Index) and MSE (Mean Squared Error).
-SSIM (range: -1 to 1) assesses similarity based on luminance, contrast, and structure, while MSE quantifies pixel-wise differences, with lower values indicating higher similarity.
-While SSIM aligns with human perception, PSNR and MSE focus on absolute pixel differences.
-(a) Comparison of models, trained on GSS-RetVein and OCTA and tested on ROSE.
-(b) Comparison of models, trained on ROSE and tested on GSS-RetVein and OCTA.
-(c) Comparison of models, trained on GSS-RetVein or OCTA and tested on each other.
-Figure 6: Conducting mixed testing on different models using mixed datasets. ($\times$100%)
-5 Experimental Result
+As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, X-GAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation.
+As shown in Table 2, comparison between CycleGAN [35] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%.
+CycleGAN (ours) consistently outperforms CycleGAN, achieving higher SSIM and lower MSE across all datasets.
+The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
+Table 2: Comparison Vessel Structure Refinement with Real OCTA ($\times$100%)
 Model Dataset
-GSS-RetVein OCTA-500 ROSE
+UTSW OCTA-500 ROSE
 6mm 3mm 6mm 3mm 3mm
@@
 CycleGAN (ours) 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
-Table 2: Comparison vessel structure refinement images with real OCTA images. ($\times$100%)
-5.1 Comparative Experiment
-As presented in Tab.
-1, experimental findings demonstrate that while models such as MedSAM, U-Net++, and YOLOv11-x exhibit competitive performance, X-GAN consistently outperforms all competing models across varying dataset resolutions, achieving the highest scores and excelling in retinal vessel segmentation, where both of the evaluation matrixs are even close to 100%, highlighting its superior segmentation accuracy and underscoring its robustness in fine-grained vascular structures segmentation.
-As shown in Fig.
-6(a), given the comparable sizes of GSS-RetVein and OCTA-500 (retinal vessel subset) and the smaller size of ROSE, we trained models on the first two and tested on the latter.
-X-GAN consistently achieved the highest IoU and Dice scores on ROSE, outperforming S2VNet, CauSSL, and MedSAM.
-To further assess generalization, we trained models on ROSE and tested them on GSS-RetVein and OCTA-500 (Fig.
-6(b)), where X-GAN again achieved the highest scores, followed by MedSAM and CauSSL.
-Finally, Fig.
-6(c) evaluates models across different scanning resolutions, with X-GAN consistently leading in segmentation accuracy.
-While MedSAM and UniverSeg performed well, U-Net and its variants showed strong but relatively lower performance.
-These results highlight GSS-RetVein’s high quality and X-GAN’s superior segmentation capability.
-Note: the highest indicator is shown in bold, while the second highest is marked with an underline. We do not consider X-GAN (MAX) in all Tabs. 1, 6 and 3.
-5.2 Ablation Experiment
+Human Evaluation: the medical team verified the corrected images.
+Luckily, since this study only aimed to segment only the main blood vessels, the small size of the capillaries caused errors in radius generation, which affected the accuracy of the indicators.
+5.2 Ablation Eperiment
 For baseline GAN model, GAN [34] and CycleGAN [35] are selected.
@@
 For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy.
-As shown in Tab.
-3, since main blood vessels make up a small proportion, the value calculated is also relatively low.
-When the minimum radius ratio is found, like X-GAN (MAx), aligning with biostatistical properties [22, 12, 5, 43, 7]—segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
+As shown in Table 3, since main blood vessels make up a small proportion, the value calculated is also relatively low, aligning with biostatistical properties [30, 31, 33, 32]—segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
+Table 3: Ablation Experiment of X-GAN ($\times$100%)
 Model Dataset
-GSS-RetVein
+UTSW
 6mm 3mm
@@
 X-GAN 99.41 99.71 98.66 99.33
-X-GAN (MAX) 99.89 99.94 99.66 99.83
-Table 3: Ablation experiment of X-GAN. ($\times$100%)
 As shown in Fig.
-7(a), the optimal choice of $R_{min}$ (0.2) enables X-GAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%.
+3(a), the optimal choice of $R_{min}$ (0.2) enables X-GAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%.
 When $R_{min}=1$, X-GAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold.
@@
 (b) The Number of Nodes $N$
-Figure 7: Ablation experiments of key parameters. ($\times$100%)
-Also, as shown in Eq.
-1 and Fig.
-7(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy.
-Given that the R values of X-GAN and X-GAN (MAX) are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
-Consequently, their curves largely overlap, with X-GAN (MAX) exhibiting a slight increase.
-5.3 Evaluation of Vessel Structure Refinement
-The comparison between CycleGAN [66] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%.
-CycleGAN (ours) consistently outperforms CycleGAN, achieving higher SSIM and lower MSE across all datasets.
-The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
-Human Evaluation: the medical team verified the corrected images.
-Luckily, since this study only aimed to segment only the main blood vessels, the small size of the capillaries caused errors in radius generation, which affected the accuracy of the indicators.
-Indicator SLSMs DFS (ours)
-GPU $\checkmark$ $\times$
-Annotation $\checkmark$ $\times$
-Training $\checkmark$ $\times$
-Explainability $\times$ $\checkmark$
-2D and 3D $\times$ $\checkmark$
-Generalization $\times$ -
-Table 4: Comparison of performance indicators between DFS (our) and SLSMs. ($\times$100%) (- indicates strong performance for a specific class.)
-(a) Segmentation of Main Retinal Vessels (2D)
-(b) Original Image (2D)
-(c) Original Image (3D)
-(d) Segmentation (2D)
-(e) Segmentation (3D)
-Figure 8: Segmentation results of X-GAN on GSS-RetVein (2D and 3D).
+Figure 3: Ablation Experiments of Key Parameters ($\times$100%)
+Also, as shown in Fig.
+3(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy.
+Given that the R values of X-GAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
 5.3 Efficiency Evaluation
-All in all, as shown in Tab.
-4, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity.
+Compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity.
 It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
-What is more, in more detail, the advantages of our approach are as follows in Tab. 5, our segmentor has demonstrated unsurpassed superiority.
+Table 4: Comparison of Performance Details between DFS (our) and SLSMs
 Model Details
@@
 DFS (ours, GPU based) $\leq$ 0.01ms 0M 0.00028 extremely low
-Table 5: Comparison of performance details between DFS (our) and SLSMs.
+Compared to the segmentor, our adjusted DFS offers unmatched advantages, as shown in Table 4.
+It requires no separate training or labeling, as DFS is a parameter-efficient algorithm.
+By eliminating manual annotation, it removes subjective bias from doctors.
+Additionally, its radius-based segmentation avoids pixel mapping errors in SLSMs, such as interference at vessel edges, inaccuracies at junctions, and mask coverage errors.
+What is more, the key issue is that both SLSMs and GAN+SLSMs frameworks are susceptible to fitting problems in segmentation tasks.
+This arises from the training limitations of the SLSMs model and the nature of the data.
+In contrast, DFS, as an algorithm rather than a model, is unaffected by this issue.
 5.4 Visualization
 As shown in Fig.
-8, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis.
+4, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis.
 In Fig.
-8(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow.
+4(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow.
 DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries.
 Extending this to 3D, we achieve precise main vessel extraction, with Fig.
-8(e) preserving vessel integrity and offering a clearer structure than the original Fig.
-8(c).
+4(c) preserving vessel integrity and offering a clearer structure than the original image (Fig.
+4(b)).
+(a) Segmentation of Main Retinal Vessels (2D)
+(b) Original Image (3D)
+(c) Segmentation (3D)
+Figure 4: Segmentation of X-GAN (2D and 3D)
 5.5 Analysis
-Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Tabs.
-1, 6 and 3).
-Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [22, 12, 5, 43, 7].
+Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1 and Table 3).
+Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [30, 31, 33, 32].
 We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results.
+Unlike GAN+SLSMs or SLSMs, X-GAN does not require additional data to avoid fitting issues.
+Its segmentor is efficient, computationally friendly, and free from pixel mapping errors associated with SLSMs segmentation.
 Studies [28, 29] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models.
-There are many examples: Faster R-CNN [48] and Mask R-CNN [16], YOLOv3 [47], YOLOv4 [4] and YOLOv5 [17], DeepSeek [14, 33] and ChatGPT [1].
-To the best of our knowledge, X-GAN is the best model currently, tailored to segment OCTA retinal main vessel.
 6 Conclusion
 In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs.
-We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation.
 Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability.
 Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
-Acknowledgements
-This research is sponsored by NASA 80NSSC22K0144, UTSW GMO No. 241213, and NIH 1R01AG083179-01.
+7 Future Work
+With access to state-of-the-art glaucoma detection equipment, we have curated a high-quality image dataset featuring precisely annotated retinal vascular images.
+This dataset is well-suited for glaucoma pathology research and computer vision.

### v4

@@
-X-GAN: A Generative AI-Powered Unsupervised Model for Main Vessel Segmentation of Glaucoma Screening
+GlaGAN: A Generative Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma
 Abstract
-Structural changes in main retinal blood vessels serve as critical biomarkers for the onset and progression of glaucoma.
-Identifying these vessels is vital for vascular modeling yet highly challenging.
-This paper proposes X-GAN, a generative AI-powered unsupervised segmentation model designed for extracting main blood vessels from Optical Coherence Tomography Angiography (OCTA) images.
-The process begins with the Space Colonization Algorithm (SCA) to rapidly generate a skeleton of vessels, featuring their radii.
-By synergistically integrating the generative adversarial network (GAN) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels.
-Based on this reconstruction, X-GAN achieves nearly 100% segmentation accuracy without relying on labeled data or high-performance computing resources.
-Experimental results confirm X-GAN’s superiority in evaluating main vessel segmentation compared to existing deep learning models.
-Code is here: https://github.com/VikiXie/SatMar8.
-Keywords:
-Medical Imaging Unsupervised Learning Glaucoma Pathological Segmentation Generative AI
+Structural changes in the main retinal blood vessels are critical biomarkers for glaucoma onset and progression.
+Identifying these vessels is essential for vascular modeling yet highly challenging.
+This paper introduces GlaGAN, an unsupervised generative AI model for segmenting main blood vessels in Optical Coherence Tomography Angiography (OCTA) images.
+The process begins with the Space Colonization Algorithm (SCA) to rapidly generate vessel skeletons, including radius estimations.
+By synergistically integrating generative adversarial networks (GANs) with biostatistical modeling of vessel radii, GlaGAN efficiently reconstructs 2D and 3D representations, achieving nearly 100% segmentation accuracy without requiring labeled data or high-performance computing resources.
+To address data scarcity, we also present GSS-RetVein, a high-definition mixed 2D/3D glaucoma retinal dataset featuring clear capillary structures.
+Designed for robustness testing, GSS-RetVein incorporates controlled noise while maintaining sharp capillary boundaries in 2D and enhancing 3D vascular reconstruction for blood flow prediction and glaucoma progression simulations.
+Experimental results demonstrate GSS-RetVein outperforms existing datasets in evaluating main vessel segmentation.
+Code and dataset are available: https://github.com/VikiXie/SatMar8.
+{IEEEkeywords}
+Glaucoma, Generative Adversarial Network, Space Colonization Algorithm, Medical Imaging, Main Blood Vessel Extraction, Image Segmentation
 1 Introduction
+[…]
 Glaucoma is a leading cause of irreversible blindness, often progressing silently until significant vision loss occurs [40, 37, 38, 42, 29].
-Recent research highlights choroidal microvasculature dropout as a potential biomarker for disease progression [42].
-Optical Coherence Tomography Angiography (OCTA) enables visualization of these vascular changes, but accurately assessing choroidal vessel density remains a challenging task [1, 17].
+Recent research suggests that choroidal microvasculature dropout may serve as a potential biomarker for disease progression [59, 60].
+While Optical Coherence Tomography Angiography (OCTA) allows detailed visualization of these vascular changes, accurately quantifying choroidal vessel density remains a major challenge [1, 17].
 Existing medical imaging analytics in glaucoma research predominantly depend on computer vision techniques [2, 46, 45, 44, 17, 37, 38, 36, 39, 41, 40, 47].
@@
 OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [1, 4].
-However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [42], and complex capillary structures.
+However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [43, 42], and complex capillary structures.
 Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [42, 29].
@@
 raising the question of whether segmentation can bypass pixel-based mapping and leverage vessel radius as a key marker.
-To address this, we propose X-GAN, an unsupervised model that integrates GANs [34] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training.
+To address this, we propose GlaGAN, an unsupervised model that integrates GANs [34] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training.
+To tackle data scarcity, we introduce GSS-RetVein, a benchmark dataset with clear capillary distribution in 2D and 3D, enhancing segmentation, 3D reconstruction, and blood flow prediction.
 In summary, the contributions of this paper are three-fold:
-  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images without annotations.
-  • We eliminate the reliance on synthetic label supervision mechanisms (SLSMs) by introducing a novel segmentation strategy based on vessel radius thresholds and depth-first search (DFS), enabling accurate main vessel extraction without any pixel-level annotations.
-  • Experimental results of X-GAN for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) current models, achieving nearly 100% segmentation accuracy.
+  • We introduce GlaGAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images.
+  • We propose the GSS-RetVein, a unique and pioneering dataset focused exclusively on retinal vascularity.
+GSS-RetVein consists of 2D and 3D imaging data, and also preserves intricate capillary details without compromising resolution, ensuring sharp blood vessel boundaries and minimal pixel-induced blurring.
+  • Experimental results of GlaGAN on GSS-RetVein, OCTA-500 [1] and ROSE [4] for retinal main vessel segmentation have demonstrate that GlaGAN outperformed state-of-the-art (SOTA) cureent models, achieving nearly 100% segmentation accuracy.
+Also, experimental results of different models on GSS-RetVein have shown the superiority of GSS-RetVein.
 2 Related Work
-2.1 Medical Imaging in Glaucoma
-Many datasets of glaucoma focus on computer vision tasks [37, 38, 36, 39, 41, 40, 43, 4, 1, 53, 54, 52], especially in medical image processing.
+2.1 Datasets for Glaucoma
+Many datasets of glaucoma focus on computer vision tasks [37, 38, 36, 39, 41, 40, 44, 4, 1, 54, 55, 53].
 For inner retinal blood vessels, datasets like LAG [44], OCTA-500 [1] and ROSE [4], provide various-sizeimages (e.g., 3mm, 6mm).
@@
 Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
-2.2 AI for Glaucoma
-Even with accessible data, labeling remains challenging due to annotation costs and inter-observer variability, introducing bias and hindering model performance [37, 38, 36, 39, 41, 40, 46].
-Generative AI can alleviate data scarcity and labeling constraints, though human oversight is often required [50].
-GANs [34, 35, 18, 24, 27, 25], as unsupervised models, address both issues by generating realistic samples without manual labels.
-In this framework, segmentation is typically achieved by coupling the GAN with a dedicated segmentor $S$, which can be implemented using pre-trained CNNs [45, 44, 43], U-Net variants [10, 7, 46, 12], ViT [22], Mask R-CNN [9], or MedSAM [8], enabling end-to-end vessel synthesis and segmentation.
+Figure 1: The Architecture of GlaGAN
+2.2 Computer Vision for Glaucoma
+Even when data is accessible, challenges in labeling and the subjective variability among doctors may introduce biases or hinder optimal model fitting [37, 38, 36, 39, 41, 40, 47].
+Generative AI can compensate for data deficiencies and labeling issues, though it often requires human verification [51] and generated data also face challenges in labeling.
+However, GAN [34, 35, 19, 25, 28, 26] is an unsupervised model which can solve both issues: data scarity and labeling.
+GAN-based segmentation methods are generally designed in conjunction with a dedicated segmentation network.
+For the segmentor part $S$, many pre-trained models like CNNs [46, 45, 44], U-Net [10, 7, 47, 12], ViT [23], Mask R-CNN [9] and MedSAM [8] can be used for final segmentation.
+Most recent studies from CVPR [18, 19, 21, 22], IEEE Transactions on Medical Imaging [17, 26, 27, 28] and other top conferences or journals [20, 24, 25] also follow this way.
 2.3 Motivation
-In traditional models such as GAN+SLSM [17, 25, 20, 26, 27, 18] and standalone SLSMs [10, 46, 12, 13, 44, 45, 43, 9, 8, 2, 3, 14, 7, 15], pixel-level segmentations are prone to inter-observer variability [40, 1, 2].
-However, in OCTA, retinal vessels have consistent visual traits, uniform grayscale gradients and path-like shapes, making them suitable for coordinate-based modeling with associated widths.
-We adopt the Space Colonization Algorithm (SCA) [49] to represent vessels as structured graphs, enhancing topological coherence and connectivity.
-To bridge the domain gap between initial vessel maps and real OCTA data, we apply a CycleGAN-based refinement module [35], aligning generated structures with realistic image characteristics while preserving topology.
-Unlike prior methods that segment post-rendered images, we introduce a DFS-based segmentation approach [51] applied directly to graph-structured (coordinates, radius) data.
-This intercepts the generation process to ensure anatomical fidelity, governed by a radius threshold $R_{min}$ derived from biostatistical vessel metrics [30, 31, 33, 32].
-3 Methodology
-Figure 1: The architecture of X-GAN where red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
-The complete framework of X-GAN is illustrated in Fig. 1 and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
-3.1 Vessel Structure Refinement Module
+For GAN+SLSM [17, 26, 21, 27, 28, 19], or SLSMs [10, 47, 12, 13, 45, 46, 44, 9, 8, 2, 3, 14, 7, 15], pixel-level representations can be subjective, as different doctors may annotate the details of the same object differently [40, 1, 2].
+In OCTA imaging, retinal blood vessels exhibit a uniform color change and have a shape closely resembling a path.
+They can be represented using coordinates along with road width.
+Space Colonization Algorithm (SCA) [50] just meets this requirement.
 Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity.
-The graph consists of the following components, shown in Equation 1:
+The graph consists of the following components, shown in Eq.
+1:
 $V=\left\{(x_{i},y_{i},z_{i}|r_{i})\right\}_{i=1}^{N}$ (1)
@@
 To mitigate this gap, we employ a tuned GAN based on CycleGAN [35] as a structural refinement module, aligning vessel representations with real OCTA distributions while preserving vascular topology.
+We introduce the Depth-First Search (DFS, a nineteenth century algorithm)-based segmentation approach [52], applying DFS directly to structured (coordinates, radius) data before image rendering.
+Unlike GAN+SLSM [17, 26, 21, 27, 28, 19], which segments generated images mixed with real data, our method intercepts the GAN generation process, ensuring segmentation aligns with biostatistical vessel radii.
+The DFS threshold, $R_{min}$, is defined as the minimum radius ratio of main vessels, following established biostatistical principles [29, 30, 31, 33, 32].
+Also, we introduce GSS-RetVein, a dataset capturing rich details of both main vessels and capillaries.
+The dense capillaries of our dataset introduce noise, effectively testing model robustness for SLSMs.
+This detail aids vascular reconstruction, blood flow prediction, and glaucoma progression simulation by enhancing edge and endpoint learning [1, 2, 3].
+Inspired by Harvard glaucoma dataset series [37, 38, 36, 39, 41, 40], we integrate fairness measures to minimize bias and ensure equitable model performance.
+3 Methodology
+The complete framework of GlaGAN is illustrated in Fig.
+1 where red indicates that GlaGAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
+and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
+3.1 Vessel Structure Refinement Module
 Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$.
 To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$.
-This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Equation 2:
+This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Eq.
+2:
 $L_{\text{seg}}=\mathbb{E}_{X_{r}}\left[\|S(X_{r}^{\prime})-S(X_{r})\|_{1}\right]$ (2)
@@
 $L_{\text{total}}$ $=L_{\text{GAN}}+\lambda_{\text{GP}}L_{\text{GP}}+\lambda_{\text{seg}}L_{\text{seg}}$ (5)
-where Equation 3 is the Wasserstein Adversarial Loss, Equation 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Equation 5 is the Structural Consistency via Segmentation Loss.
+where Eq.
+3 is the Wasserstein Adversarial Loss, Eq.
+4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Eq.
+5 is the Structural Consistency via Segmentation Loss.
 Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
@@
 First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity.
-A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Equation 6:
+A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Eq.
+6:
 $w(e_{ij})=r_{i}$ (6)
@@
 To extract primary vessels, we apply radius thresholding and DFS traversal.
-For radius filtering, we define the ratio of the smallest main vessel radius to the largest main vessel radius $R_{min}$ and retain only Equation 7:
+For radius filtering, we define a minimum main vessel radius $R_{min}$ and retain only Eq.
+7:
 $V_{\text{main}}=\left\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\min}*r_{max})\right\}$ (7)
-For DFS traversal, it consists of three steps: (1) select the optic disc region as the root node; (2) recursively traverse the largest connected component using DFS; (3) stop when no further vessel segments satisfy the radius constraint.
+For DFS traversal, it consists of three steps:
+  • select the optic disc region as the root node
+  • recursively traverse the largest connected component using DFS
+  • stop when no further vessel segments satisfy the radius constraint
 when all three steps finished, construct final extracted vessel structure $G_{main}$.
 This approach effectively retains only the primary vascular structure, removing small capillaries and noise.
+This method can be summarized as the following Algorithm 1.
+Algorithm 1 Primary Vessel Extraction via DFS
+0:  Vessel Structure $V=\{(x_{i},y_{i},z_{i}\mid r_{i})\}_{i=1}^{N}$
+0:  Extracted Primary Vessel Graph $G_{\text{main}}$
+1:  Step 1: Vessel Graph Construction
+2:  for all vessel point $(x_{i},y_{i},z_{i}\mid r_{i})\in V$ do
+3:     Identify Neighboring Vessel Points to Form Edge Set $E$
+4:     Assign Edge Weight $w(e_{ij})=r_{i}$
+5:  end for
+6:  Step 2: Vessel Filtering
+7:  Apply Radius Threshold $R_{\text{min}}$
+8:  $V_{\text{main}}=\{(x_{i},y_{i},z_{i}\mid r_{i})\mid r_{i}\geq R_{\text{min}}\cdot r_{\max}\}$
+9:  Remove Weakly Connected Components
+10:  Step 3: DFS-Based Traversal
+11:  Initialize DFS from Optic Disc Region
+12:  for all point $(x_{i},y_{i},z_{i}\mid r_{i})\in V_{\text{main}}$ do
+13:     Recursively Explore Vessel Network via DFS
+14:     Preferentially Expand along Thicker Vessels
+15:  end for
+16:  Step 4: Output
+17:  return  $G_{\text{main}}$
+The key aspect of Algorithm 1 is the threshold $R_{min}$.
+Instead of the final image, we analyze intermediate outputs $S(X_{r})$, CSV data (Fig.
+2(e)), and GAN-derived radius statistics (Fig.
+2(d)).
+For a sample image (Fig.
+2(a)), we computed a Mean (85.31) and Standard Deviation (88.67), determining $R_{min}$ as 0.1968 $r_{max}$.
+Across datasets, values were 0.2011 (GSS-RetVein), 0.2102 (OCTA-500), and 0.1982 (ROSE), standardized to 0.2 for consistency.
+Ophthalmology studies [29, 30, 31, 33, 32] confirm 18-26% capillary coverage in OCTA images, validating this threshold.
+(a) GSS-RetVein
+(b) 3D Visualization
+(c) Statistical Distribution
+(d) Radius Statistics
+(e) Vascular Data
+Figure 2: 3D Visualization of the Pixel Distribution of One Sample, and its Gaussian Statistical Distribution
 4 Dataset and Implementation
 4.1 Dataset
+4.1.1 GSS-RetVein
+GSS-RetVein is a unique and pioneering dataset focused exclusively on retinal vascularity.
+The dataset features an exceptionally clear distribution of capillaries.
+It comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race.
+We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female.
+Racial distribution reflects local glaucoma clinic demographics: 80.4% white, 14.6% Black, with Asians and other minority groups comprising the remainder.
+Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
+GSS-RetVein’s original images come from the latest Intalight device [^1] https://intalight.com/, offering superior detail compared to OPTOVUE [^2] https://www.visionix.com/ (Fig.
+3).
+As shown in Fig.
+3(a) and Fig.
+3(c), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy.
+In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [3].
+For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [2], making Intalight data ideal for both tasks.
+(a) Intalight
+(b) OPTOVUE
+(c) Intalight
+(d) OPTOVUE
+Figure 3: Comparison of Data Collected by Different Devices for the Same Patient
+For GSS-RetVein, we eliminate red and green baselines from Intalight device images through a multi-step process.
+First, we apply a dark channel algorithm [5] to remove light artifacts, followed by unsharp masking [6] to enhance high-frequency details.
+The final images (Fig.
+4(a)) undergo manual verification for quality assurance.
+As shown in Fig.
+4, while main vessels remain visible across datasets, magnification reveals that OCTA-500 [1] and ROSE [4] blur capillaries and vessel ends.
+Even the unprocessed baseline image (Fig.
+LABEL:cid) offers superior vessel clarity compared to other datasets (Fig.
+4(c), Fig.
+4(b)).
+Among them, Fig.
+3(a) and Fig.
+3(b) are images of deep angio, and Fig.
+3(c) and Fig.
+3(d) are images of superficial angio.
+They all are images of inner retina vessel.
+(a) GSS-RetVein
+(b) OCTA-500
+(c) ROSE
+Figure 4: Comparison of GSS-RetVein with Other Datasets
+For the representation of 3D imaging of GSS-RetVein, as shown in Fig.
+5, capillaries stand out from the thicker main blood vessels by attaching to them and narrowing at their ends (Fig.
+5(c)).
+The elaborate architecture of blood vessels and the intertwining capillaries contribute to an extremely intricate vascular layout.
+3D imaging part of GSS-RetVein is formatted as a JSON file, which includes 3D X, Y, and Z coordinates and the corresponding radius pixel distribution values.
+For annotation (2D, image format: PNG), our team includes two glaucoma specialists who handle labeling, verification, and manual evaluation, utilizing the Labelme [^3] https://github.com/wkentaro/labelme.
+(a) Top View
+(b) Front View
+(c) Detail
+Figure 5: 3D Imaging of Retinal Vessel from Various Angles
+In general, GSS-RetVein is a dataset that combines 2D and 3D images with extremely high image resolution, and is particularly dedicated to retinal vascular segmentation.
+Table 1: Comparison GlaGAN With Other Models on GSS-RetVein, OCTA-500 and ROSE ($\times$100%)
+Model Dataset
+GSS-RetVein OCTA-500 ROSE
+6mm 3mm 6mm 3mm 3mm
+IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
+U-Net 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
+U-Net++ 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
+Attention U-Net 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
+Mask R-CNN 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
+YOLOv11-x 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
+MedSAM 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
+CauSSL 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
+UniverSeg 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
+S2VNet 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
+Tyche 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
+GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
+Table 2: Conducting Mixed Validation on Different Models using Mixed Datasets ($\times$100%)
+Model Dataset
+ROSE GSS-RetVein OCTA-500 GSS-RetVein (Test) OCTA-500 (Test)
+GSS-RetVein OCTA-500 ROSE OCTA-500 (Train) GSS-RetVein (Train)
+3mm 6mm 3mm 6mm 3mm
+IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
+U-Net 98.34 99.16 98.32 99.15 93.89 96.85 94.23 97.03 96.21 98.07 95.23 97.56 96.98 98.47 96.81 98.38
+U-Net++ 98.99 99.49 98.78 99.39 94.08 96.95 94.46 97.15 95.73 97.82 97.01 98.51 97.59 98.78 96.93 98.44
+Attention U-Net 98.54 99.26 98.23 99.11 94.07 96.94 94.52 97.18 96.99 98.47 96.09 98.01 97.21 98.59 96.78 98.36
+Mask R-CNN 96.37 98.15 96.31 98.12 90.73 95.14 91.89 95.77 94.23 97.03 94.11 96.97 96.68 98.31 95.96 97.94
+YOLOv11-x 98.88 99.44 98.72 99.36 94.11 96.97 94.49 97.17 97.32 98.64 96.54 98.24 96.12 98.02 96.63 98.29
+MedSAM 99.04 99.52 98.98 99.49 95.01 97.41 95.42 97.66 97.54 98.75 95.87 97.89 98.23 99.11 97.46 98.72
+CauSSL 96.83 98.38 95.89 97.90 96.98 98.47 97.83 98.88 99.12 99.56 98.81 99.40 94.87 97.33 95.34 97.62
+UniverSeg 98.96 99.48 98.56 99.28 94.23 97.03 94.59 97.24 96.81 98.38 96.32 98.16 97.91 98.96 98.12 99.05
+S2VNet 99.23 99.61 98.79 99.39 93.89 96.85 95.22 97.55 96.95 98.44 95.83 97.87 97.87 98.95 97.87 98.95
+Tyche 98.78 99.38 98.43 99.21 94.22 97.02 94.98 97.46 96.88 98.41 96.11 98.02 97.84 98.94 97.68 98.87
+GlaGAN 99.55 99.78 99.21 99.60 96.98 98.46 97.69 98.83 98.87 99.43 98.79 99.39 99.58 99.79 99.41 99.70
+Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki.
+Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
 4.1.2 OCTA-500
@@
 ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
-(a) UTSW Data
-(b) OCTA-500
-(c) ROSE
-Figure 2: Comparison of UTSW with Other Datasets
-4.1.3 UTSW
-our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki.
-Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
-As show in Fig.
-2, original images come from the latest Intalight device [^1] https://intalight.com/, which is currently the most advanced equipment.
-As shown in Fig.
-2(a), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy.
-UTSW comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race.
-We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female.
-Racial distribution reflects local glaucoma clinic demographics: 80.4% white, 14.6% Black, with Asians and other minority groups comprising the remainder.
-Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
-In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [3].
-For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [2], making Intalight data ideal for both tasks.
-Our team includes two attending glaucoma physicians who are responsible for the detailed labeling of the main retinal blood vessels using the tool Labelme [^2] https://github.com/wkentaro/labelme.
-4.2 Implementation
+4.2 Implement
+Table 3: Comparison Vessel Structure Refinement Images with Real OCTA Images ($\times$100%)
+Model Dataset
+GSS-RetVein OCTA-500 ROSE
+6mm 3mm 6mm 3mm 3mm
+SSIM MSE SSIM MSE SSIM MSE SSIM MSE SSIM MSE
+GAN 91.76 5.69 90.39 4.43 89.69 9.85 91.06 7.36 87.84 8.73
+CycleGAN 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
+GAN 93.62 3.29 93.87 3.79 91.28 6.98 93.19 5.54 91.73 4.18
+CycleGAN 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
 4.2.1 Hyperparameter
-For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70).
+for GlaGAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70).
 The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$.
-X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation.
-Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$) and flipping.
-4.2.2 Evaluation Matrix
+GlaGAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation.
+Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
+4.2.2 Hardware
+the hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
+4.2.3 Evaluation Metric
 we use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth.
@@
 While SSIM aligns with human perception, MSE focus on absolute pixel differences.
-4.2.3 Experimental Setup
-We selected the following deep learning models as reference benchmarks: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [13], UniverSeg [14], S2VNet [15] and Tyche [16].
-These models are leaders in the fields of image segmentation and medical image processing.
-For their parameter adjustment, we use Optuna [^3] https://optuna.org/ to optimize their learning green, decay and other hyperparameters.
-For the training and testing of the model, we use ten cross-validation methods.
-The evaluation indicators are finally averaged.
-4.2.4 Hardware
-The hardware for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
-5 Experimental Result
-5.1 Comparative Experiment
-Table 1: Comparison X-GAN with Other Models on Different Datasets ($\times$100%)
+Table 4: Ablation Experiment of GlaGAN ($\times$100%)
 Model Dataset
-UTSW OCTA-500 ROSE
+GSS-RetVein OCTA-500 ROSE
 6mm 3mm 6mm 3mm 3mm
 IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
-U-Net [10] 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
-U-Net++ [7] 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
-Attention U-Net [12] 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
-Mask R-CNN [9] 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
-YOLOv11-x [11] 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
-MedSAM [8] 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
-CauSSL [13] 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
-UniverSeg [14] 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
-S2VNet [15] 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
-Tyche [16] 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
-X-GAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
-As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, X-GAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation.
+GAN U-Net 92.89 96.31 92.44 96.07 93.31 96.56 92.84 96.31 93.37 96.59
+GAN U-Net++ 93.62 96.70 92.87 96.30 93.72 96.73 93.03 96.36 92.84 96.31
+GAN Attention U-Net 93.56 96.67 93.12 96.44 93.65 96.70 92.85 96.31 93.42 96.61
+GAN MedSAM 93.73 96.76 92.76 96.24 92.83 96.30 92.47 96.10 93.76 96.75
+GAN DFS 95.81 97.86 95.24 97.56 94.83 97.33 94.87 97.35 94.73 97.27
+CycleGAN U-Net 93.44 96.61 93.42 96.60 94.73 97.26 93.97 96.89 93.76 96.75
+CycleGAN U-Net++ 93.72 96.76 93.62 96.70 94.58 97.18 94.36 97.05 93.19 96.46
+CycleGAN Attention U-Net 93.79 96.80 93.45 96.61 94.59 97.18 94.29 97.01 94.03 96.89
+CycleGAN MedSAM 94.02 96.92 94.37 97.10 93.83 96.72 94.37 97.19 94.18 96.96
+CycleGAN DFS 96.23 98.08 96.42 98.18 96.68 98.33 96.27 98.13 95.83 97.84
+CycleGAN U-Net 94.56 97.20 94.52 97.18 94.86 97.31 94.79 97.27 94.19 96.94
+CycleGAN U-Net++ 94.98 97.43 94.36 97.10 94.83 97.29 95.09 97.41 94.27 96.98
+CycleGAN Attention U-Net 95.06 97.47 95.12 97.50 95.07 97.39 94.97 97.33 94.39 97.04
+CycleGAN MedSAM 95.33 97.61 95.21 97.55 95.36 97.55 95.11 97.41 94.82 97.27
+GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
+5 Experimental Result
+5.1 Experimental Setup
+We selected many SOTA models for experimental verification, which are: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [13], UniverSeg [14], S2VNet [15], Tyche [16], GAN [34] and CycleGAN [35].
+Among them, we used tool Optuna [^4] https://optuna.org/ to optimize their parameters.
+All final experimental results are the average of ten cross-validation times.
+* means models’ hyperparameters have been optimized by us.
+5.2 Comparative Experiment
+As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, GlaGAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation.
+As shown in Table 2, the second row is the test data set, and the third row is the training data set.
+Models trained on GSS-RetVein and OCTA-500 were tested on ROSE, where GlaGAN achieved the highest IoU and Dice scores, outperforming S2VNet, CauSSL, and MedSAM.
+For generalization, training on ROSE and testing on GSS-RetVein and OCTA-500 confirmed GlaGAN’s superiority, followed by MedSAM and CauSSL.
+Across different scanning resolutions, GlaGAN consistently led in segmentation accuracy, with MedSAM and UniverSeg performing well, while U-Net variants lagged.
+These results affirm GSS-RetVein’s quality and GlaGAN’s segmentation excellence.
+5.3 Ablation Experiment
+For baseline GAN model, GAN and CycleGAN are selected.
+For baseline segmentor, U-Net, U-Net++, Attention U-Net and MedSAM are selected.
+For GlaGAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy.
+As shown in Table 4, since main blood vessels make up a small proportion, the value calculated is also relatively low.
+When the minimum radius ratio is found, like GlaGAN, aligning with biostatistical properties [29, 30, 31, 33, 32], segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
+Table 5: Ablation Experiment of Hyperparameters ($\times$100%)
+Hyperparameter Dataset
+GSS-RetVein OCTA-500 ROSE
+6mm 3mm 6mm 3mm 3mm
+IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
+Rmin 0.00 train 1.03 2.04 1.00 1.98 1.31 2.59 1.22 2.41 1.13 2.23
+test 0.75 1.49 0.87 1.73 0.91 1.80 0.77 1.52 0.93 1.84
+0.10 train 41.32 58.47 41.98 59.12 43.09 60.23 45.57 62.59 44.29 61.38
+test 40.29 57.42 42.75 59.89 42.31 59.45 45.06 62.13 43.97 61.08
+0.20 train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
+test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
+0.30 train 87.15 93.13 86.29 92.64 86.04 92.49 85.74 92.30 84.37 91.52
+test 84.17 91.40 83.96 91.27 84.09 91.35 84.02 91.31 83.98 91.28
+0.40 train 57.92 73.34 56.44 72.15 55.86 71.65 55.45 71.28 57.09 72.69
+test 58.73 73.98 57.85 73.29 57.29 72.85 56.93 72.53 57.24 72.80
+0.50 train 55.72 71.56 56.98 72.60 54.79 70.78 56.93 72.53 57.04 72.65
+test 54.17 70.25 53.83 69.99 52.62 68.97 53.71 69.89 53.83 69.99
+0.60 train 40.26 57.40 41.27 58.42 40.39 57.53 41.03 58.17 39.27 56.39
+test 35.62 52.51 38.92 56.03 37.46 54.50 37.08 54.10 35.74 52.63
+0.70 train 35.36 52.25 34.72 51.55 33.91 50.64 37.06 54.08 35.94 52.85
+test 32.24 48.76 31.83 48.27 32.64 49.21 33.39 50.06 34.07 50.83
+0.80 train 17.83 30.26 19.25 32.29 18.83 31.69 17.79 30.21 16.68 28.59
+test 26.37 41.73 25.73 40.93 25.49 40.61 25.13 40.15 24.80 39.73
+0.90 train 9.85 17.94 10.32 18.71 9.89 17.99 11.27 20.26 9.91 18.03
+test 12.24 21.82 13.06 23.10 12.94 22.92 11.98 21.39 10.57 19.12
+1.00 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Node 0 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+5000 train 6.34 11.92 7.81 14.49 6.28 11.82 6.43 12.09 7.02 13.12
+test 6.06 11.43 5.86 11.07 5.93 11.19 6.18 11.64 6.58 12.35
+10000 train 17.82 30.24 18.92 31.81 17.96 30.44 19.06 32.02 17.48 29.75
+test 20.59 34.15 21.76 35.73 22.84 37.19 21.96 36.00 20.98 34.69
+15000 train 20.58 34.14 21.96 36.00 20.84 34.50 21.37 35.21 20.71 34.29
+test 28.73 44.63 29.64 45.71 30.74 47.04 28.93 44.87 29.74 45.83
+20000 train 41.47 58.61 42.34 59.49 41.65 58.80 42.07 59.18 44.03 61.14
+test 42.06 59.17 41.11 58.26 41.93 59.04 40.62 57.77 41.98 59.12
+25000 train 49.38 66.12 52.14 68.53 50.39 66.96 52.97 69.26 51.07 67.60
+test 52.01 68.41 51.49 67.98 51.76 68.21 52.44 68.79 53.06 69.35
+30000 train 59.23 74.39 59.17 74.34 58.48 73.78 57.99 73.40 58.49 73.79
+test 61.25 75.97 60.84 75.64 59.73 74.78 61.79 76.37 60.83 75.63
+35000 train 62.33 76.79 63.34 77.56 62.15 76.64 61.96 76.49 62.15 76.64
+test 62.51 76.96 64.08 78.12 63.24 77.47 62.44 76.89 63.19 77.43
+40000 train 62.74 77.13 64.27 78.25 63.74 77.88 63.96 78.06 63.49 77.64
+test 62.59 77.03 64.13 78.16 63.96 78.06 61.84 76.39 63.97 78.07
+45000 train 63.37 77.57 65.03 78.81 65.17 79.57 65.03 78.81 65.93 79.41
+test 65.72 79.30 66.34 79.77 64.44 78.39 65.39 78.95 66.67 80.00
+50000 train 66.53 79.89 66.02 79.53 65.52 79.14 65.74 79.31 66.03 79.54
+test 69.34 81.89 68.18 81.06 69.01 81.65 66.95 80.19 68.39 81.24
+50000 train 71.24 83.20 72.47 84.02 69.08 81.71 68.79 81.52 69.28 81.83
+test 73.69 84.85 72.61 84.12 74.17 85.15 75.36 85.94 74.13 85.10
+55000 train 75.48 85.99 76.24 86.54 76.14 86.47 75.83 86.25 74.19 85.17
+test 76.02 86.38 75.47 85.98 74.45 85.34 73.99 85.05 74.73 85.57
+60000 train 79.28 88.47 78.96 88.23 79.48 88.61 77.20 87.13 75.96 86.35
+test 78.49 87.94 77.85 87.52 77.02 87.03 79.36 88.52 75.89 86.29
+65000 train 85.27 92.05 86.09 92.52 85.78 92.29 85.96 92.42 86.06 92.50
+test 84.18 91.40 84.07 91.33 83.33 90.89 82.39 90.37 83.09 90.76
+70000 train 97.46 98.71 97.28 98.62 97.08 98.52 96.96 98.46 97.07 98.51
+test 98.32 99.15 98.29 99.14 99.03 99.51 99.01 99.50 98.76 99.37
+…… train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
+test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
+As shown in Table 5, the optimal choice of $R_{min}$ (0.2) enables GlaGAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%.
+When $R_{min}=1$, GlaGAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold.
+When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
+Also, as shown in Fig.
+ref1 and Table 5, a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy.
+Given that the R values of GlaGAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
+Consequently, their curves largely overlap, with GlaGAN exhibiting a slight increase.
+5.4 Evaluation of Vessel Structure Refinement
 As shown in Table 3, comparison between CycleGAN [35] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%.
@@
 The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
-Table 2: Comparison Vessel Structure Refinement with Real OCTA ($\times$100%)
-Model Dataset
-UTSW OCTA-500 ROSE
-6mm 3mm 6mm 3mm 3mm
-SSIM MSE SSIM MSE SSIM MSE SSIM MSE SSIM MSE
-CycleGAN [35] 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
-CycleGAN (ours) 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
-Human Evaluation: the medical team verified the corrected images.
-Luckily, since this study only aimed to segment only the main blood vessels, the small size of the capillaries caused errors in radius generation, which affected the accuracy of the indicators.
-5.2 Ablation Eperiment
-For baseline GAN model, GAN [34] and CycleGAN [35] are selected.
-For baseline segmentor, U-Net [10], U-Net++ [7], Attention U-Net [12] and MedSAM [8] are selected.
-For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy.
-As shown in Table 3, since main blood vessels make up a small proportion, the value calculated is also relatively low, aligning with biostatistical properties [30, 31, 33, 32]—segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
-Table 3: Ablation Experiment of X-GAN ($\times$100%)
-Model Dataset
-UTSW
-6mm 3mm
-IoU Dice IoU Dice
-GAN [34] + U-Net [10] 92.89 96.31 92.44 96.07
-GAN [34] + U-Net++ [7] 93.62 96.70 92.87 96.30
-GAN [34] + Attention U-Net [12] 93.56 96.67 93.12 96.44
-GAN [34] + MedSAM [8] 93.73 96.76 92.76 96.24
-GAN [34] + DFS ($R_{min}$=0.2) 95.81 97.86 95.24 97.56
-CycleGAN [35] + U-Net [10] 93.44 96.61 93.42 96.60
-CycleGAN [35] + U-Net++ [7] 93.72 96.76 93.62 96.70
-CycleGAN [35] + Attention U-Net [12] 93.79 96.80 93.45 96.61
-CycleGAN [35] + MedSAM [8] 94.02 96.92 94.37 97.10
-CycleGAN [35] + DFS ($R_{min}$=0.2) 96.23 98.08 96.42 98.18
-CycleGAN (ours) + U-Net [10] 94.56 97.20 94.52 97.18
-CycleGAN (ours) + U-Net++ [7] 94.98 97.43 94.36 97.10
-CycleGAN (ours) + Attention U-Net [12] 95.06 97.47 95.12 97.50
-CycleGAN (ours) + MedSAM [8] 95.33 97.61 95.21 97.55
-X-GAN 99.41 99.71 98.66 99.33
-As shown in Fig.
-3(a), the optimal choice of $R_{min}$ (0.2) enables X-GAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%.
-When $R_{min}=1$, X-GAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold.
-When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
-(a) $R_{min}$
-(b) The Number of Nodes $N$
-Figure 3: Ablation Experiments of Key Parameters ($\times$100%)
-Also, as shown in Fig.
-3(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy.
-Given that the R values of X-GAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
 5.5 Efficiency Evaluation
-Compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity.
+As shown in Table 6, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity.
 It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
-Table 4: Comparison of Performance Details between DFS (our) and SLSMs
+Table 6: Comparison of Performance Indicators between DFS and SLSMs ($\times$100%)
+Indicator SLSMs DFS (ours)
+GPU $\checkmark$ $\times$
+Annotation $\checkmark$ $\times$
+Training $\checkmark$ $\times$
+Explainability $\times$ $\checkmark$
+2D and 3D $\times$ $\checkmark$
+Generalization $\times$ -
 Model Details
 Inference Time Parameters FLOPs GPU Utilization
-U-Net [10] 21ms 31M 98 low
-U-Net++ [7] 62ms 76M 304 high
-Attention U-Net [12] 83ms 62M 215 slightly high
-MedSAM [8] 425ms 109M 1250 extremely high
-DFS (ours) 0.018ms 0M 0.00028 -
-DFS (ours, GPU based) $\leq$ 0.01ms 0M 0.00028 extremely low
+U-Net 21ms 31M 98 low
+U-Net++ 62ms 76M 304 high
+Attention U-Net 83ms 62M 215 slightly high
+MedSAM 425ms 109M 1250 extremely high
+DFS (CPU) 0.018ms 0M 0.00028 -
+DFS (GPU) $\leq$ 0.01ms 0M 0.00028 extremely low
+Table 7: Comparison of Performance Details between DFS and SLSMs
 Compared to the segmentor, our adjusted DFS offers unmatched advantages, as shown in Table 7.
@@
 As shown in Fig.
-4, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis.
+6, 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis.
 In Fig.
-4(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow.
+6(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow.
 DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries.
 Extending this to 3D, we achieve precise main vessel extraction, with Fig.
-4(c) preserving vessel integrity and offering a clearer structure than the original image (Fig.
-4(b)).
+6(e) preserving vessel integrity and offering a clearer structure than the original image (Fig.
+6(c)).
 (a) Segmentation of Main Retinal Vessels (2D)
-(b) Original Image (3D)
-(c) Segmentation (3D)
-Figure 4: Segmentation of X-GAN (2D and 3D)
+(b) Original Image (2D)
+(c) Original Image (3D)
+(d) Segmentation (2D)
+(e) Segmentation (3D)
+Figure 6: Segmentation Results of GlaGAN on GSS-RetVein (2D and 3D)
 5.7 Analysis
-Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1 and Table 3).
-Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [30, 31, 33, 32].
-We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results.
-Unlike GAN+SLSMs or SLSMs, X-GAN does not require additional data to avoid fitting issues.
+Optimal selection of $R_{min}$ values allows GlaGAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1, Table 2, Table 4).
+Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [29, 30, 31, 33, 32].
+We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted GlaGAN to achieve amazing segmentation results.
+Unlike GAN+SLSMs or SLSMs, GlaGAN does not require additional data to avoid fitting issues.
 Its segmentor is efficient, computationally friendly, and free from pixel mapping errors associated with SLSMs segmentation.
-Studies [28, 29] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models.
+To the best of our knowledge, GlaGAN is the best model currently, tailored to segment OCTA retinal main vessel.
 6 Conclusion
-In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs.
-Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability.
+In this paper, we propose GlaGAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs.
+We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation.
+Extensive experiments demonstrate that GlaGAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability.
 Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
-7 Future Work
-With access to state-of-the-art glaucoma detection equipment, we have curated a high-quality image dataset featuring precisely annotated retinal vascular images.
-This dataset is well-suited for glaucoma pathology research and computer vision.
+Acknowledgment
+This research is partially supported by NASA under grant 80NSSC22K0144, UTSW GMO No. 241213, and NIH grant 1R01AG083179-01.

### v5

@@
-GlaGAN: A Generative Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma
+X-GAN: A Generative AI-Powered Unsupervised Model for Main Vessel Segmentation of Glaucoma Screening
 Abstract
-Structural changes in the main retinal blood vessels are critical biomarkers for glaucoma onset and progression.
-Identifying these vessels is essential for vascular modeling yet highly challenging.
-This paper introduces GlaGAN, an unsupervised generative AI model for segmenting main blood vessels in Optical Coherence Tomography Angiography (OCTA) images.
-The process begins with the Space Colonization Algorithm (SCA) to rapidly generate vessel skeletons, including radius estimations.
-By synergistically integrating generative adversarial networks (GANs) with biostatistical modeling of vessel radii, GlaGAN efficiently reconstructs 2D and 3D representations, achieving nearly 100% segmentation accuracy without requiring labeled data or high-performance computing resources.
-To address data scarcity, we also present GSS-RetVein, a high-definition mixed 2D/3D glaucoma retinal dataset featuring clear capillary structures.
-Designed for robustness testing, GSS-RetVein incorporates controlled noise while maintaining sharp capillary boundaries in 2D and enhancing 3D vascular reconstruction for blood flow prediction and glaucoma progression simulations.
-Experimental results demonstrate GSS-RetVein outperforms existing datasets in evaluating main vessel segmentation.
-Code and dataset are available: https://github.com/VikiXie/SatMar8.
-{IEEEkeywords}
-Glaucoma, Generative Adversarial Network, Space Colonization Algorithm, Medical Imaging, Main Blood Vessel Extraction, Image Segmentation
+Structural changes in main retinal blood vessels serve as critical biomarkers for the onset and progression of glaucoma.
+Identifying these vessels is vital for vascular modeling yet highly challenging.
+This paper proposes X-GAN, a generative AI-powered unsupervised segmentation model designed for extracting main blood vessels from Optical Coherence Tomography Angiography (OCTA) images.
+The process begins with the Space Colonization Algorithm (SCA) to rapidly generate a skeleton of vessels, featuring their radii.
+By synergistically integrating the generative adversarial network (GAN) with biostatistical modeling of vessel radii, X-GAN enables a fast reconstruction of both 2D and 3D representations of the vessels.
+Based on this reconstruction, X-GAN achieves nearly 100% segmentation accuracy without relying on labeled data or high-performance computing resources.
+Experimental results confirm X-GAN’s superiority in evaluating main vessel segmentation compared to existing deep learning models.
 1 Introduction
-[…]
 Glaucoma is a leading cause of irreversible blindness, often progressing silently until significant vision loss occurs [34, 41, 42, 44, 47].
-Recent research suggests that choroidal microvasculature dropout may serve as a potential biomarker for disease progression [59, 60].
-While Optical Coherence Tomography Angiography (OCTA) allows detailed visualization of these vascular changes, accurately quantifying choroidal vessel density remains a major challenge [1, 17].
-Existing medical imaging analytics in glaucoma research predominantly depend on computer vision techniques [2, 46, 45, 44, 17, 37, 38, 36, 39, 41, 40, 47].
-In recent years, supervised learning segmentation models (SLSMs) [10, 47, 12, 13, 45, 46, 44, 9, 8, 2, 3, 14, 7, 15] have become particularly prominent in this field.
+Recent research highlights choroidal microvasculature dropout as a potential biomarker for disease progression [47].
+Optical Coherence Tomography Angiography (OCTA) enables visualization of these vascular changes, but accurately assessing choroidal vessel density remains a challenging task [1, 13, 18, 20].
+Existing medical imaging analytics in glaucoma research predominantly depend on computer vision techniques [2, 20, 30, 40, 41, 42, 43, 44, 45, 46, 48, 49, 50, 51].
+In recent years, supervised learning segmentation models (SLSMs) [2, 3, 7, 8, 9, 10, 12, 14, 15, 16, 17, 48, 49, 50, 51] have become particularly prominent in this field.
 However, SLSMs may not be well-suited for detecting choroidal vessels in OCTA images.
@@
 OCTA images offer high-resolution, depth-resolved visualization of retinal and choroidal microvasculature without the need of contrast dye, making them more effective than traditional fundus imaging for detecting vascular abnormalities [1, 4].
-However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [43, 42], and complex capillary structures.
+However, segmenting choroidal vessels is difficult due to irregular patterns, intersecting pathways, dense capillary networks [47], and complex capillary structures.
 Moreover, the presence of major blood vessels, which must be excluded, further complicates the labeling process [34, 47].
@@
 raising the question of whether segmentation can bypass pixel-based mapping and leverage vessel radius as a key marker.
-To address this, we propose GlaGAN, an unsupervised model that integrates GANs [34] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training.
-To tackle data scarcity, we introduce GSS-RetVein, a benchmark dataset with clear capillary distribution in 2D and 3D, enhancing segmentation, 3D reconstruction, and blood flow prediction.
+To address this, we propose X-GAN, an unsupervised model that integrates GANs [38] with retinal vessel biostatistics for precise main vessel segmentation from OCTA images, eliminating the need for labeled data or extensive training.
 In summary, the contributions of this paper are three-fold:
-  • We introduce GlaGAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images.
-  • We propose the GSS-RetVein, a unique and pioneering dataset focused exclusively on retinal vascularity.
-GSS-RetVein consists of 2D and 3D imaging data, and also preserves intricate capillary details without compromising resolution, ensuring sharp blood vessel boundaries and minimal pixel-induced blurring.
-  • Experimental results of GlaGAN on GSS-RetVein, OCTA-500 [1] and ROSE [4] for retinal main vessel segmentation have demonstrate that GlaGAN outperformed state-of-the-art (SOTA) cureent models, achieving nearly 100% segmentation accuracy.
-Also, experimental results of different models on GSS-RetVein have shown the superiority of GSS-RetVein.
+  • We introduce X-GAN, a generative AI-powered unsupervised segmentation model tailored for extracting main blood vessels from OCTA images without annotations.
+  • We eliminate the reliance on synthetic label supervision mechanisms by introducing a novel segmentation strategy based on vessel radius thresholds and depth-first search (DFS), enabling accurate main vessel extraction without any pixel-level annotations.
+  • Experimental results of X-GAN for retinal main vessel segmentation have demonstrate that X-GAN outperformed state-of-the-art (SOTA) current models, achieving nearly 100% segmentation accuracy.
 2 Related Work
-2.1 Datasets for Glaucoma
-Many datasets of glaucoma focus on computer vision tasks [37, 38, 36, 39, 41, 40, 44, 4, 1, 54, 55, 53].
+2.1 Medical Imaging in Glaucoma
+Many datasets of glaucoma focus on computer vision tasks [1, 4, 40, 41, 42, 43, 44, 46, 48, 57, 58, 59], especially in medical image processing.
 For inner retinal blood vessels, datasets like LAG [48], OCTA-500 [1] and ROSE [4], provide various-sizeimages (e.g., 3mm, 6mm).
@@
 Some datasets [40, 41, 42, 43, 44, 46, 48, 57, 58, 59] offer fundus images, and the capillaries are nearly indistinguishable to the naked eye.
-There is a shortage of high-quality, intuitive data, such as detailed images of inner 3mm×$\times$3mm retinal vessels, and existing datasets often lack sufficient clarity.
+There is a shortage of high-quality, intuitive data, such as detailed images of inner $\times3mm$ retinal vessels, and existing datasets often lack sufficient clarity.
 Consequently, few studies focus on structural changes in major retinal blood vessels in glaucoma.
-Figure 1: The Architecture of GlaGAN
-2.2 Computer Vision for Glaucoma
-Even when data is accessible, challenges in labeling and the subjective variability among doctors may introduce biases or hinder optimal model fitting [37, 38, 36, 39, 41, 40, 47].
-Generative AI can compensate for data deficiencies and labeling issues, though it often requires human verification [51] and generated data also face challenges in labeling.
-However, GAN [34, 35, 19, 25, 28, 26] is an unsupervised model which can solve both issues: data scarity and labeling.
-GAN-based segmentation methods are generally designed in conjunction with a dedicated segmentation network.
-For the segmentor part $S$, many pre-trained models like CNNs [46, 45, 44], U-Net [10, 7, 47, 12], ViT [23], Mask R-CNN [9] and MedSAM [8] can be used for final segmentation.
-Most recent studies from CVPR [18, 19, 21, 22], IEEE Transactions on Medical Imaging [17, 26, 27, 28] and other top conferences or journals [20, 24, 25] also follow this way.
+2.2 AI for Glaucoma
+Even with accessible data, labeling remains challenging due to annotation costs and inter-observer variability, introducing bias and hindering model performance [40, 41, 42, 43, 44, 46, 51].
+Generative AI can alleviate data scarcity and labeling constraints, though human oversight is often required [55].
+GANs [21, 27, 28, 31, 38, 39], as unsupervised models, address both issues by generating realistic samples without manual labels.
+In this framework, segmentation is typically achieved by coupling the GAN with a dedicated segmentor $S$, which can be implemented using pre-trained CNNs [48, 49, 50], U-Net variants [7, 10, 12, 51], ViT [25], Mask R-CNN [9], or MedSAM [8], enabling end-to-end vessel synthesis and segmentation.
 2.3 Motivation
-For GAN+SLSM [17, 26, 21, 27, 28, 19], or SLSMs [10, 47, 12, 13, 45, 46, 44, 9, 8, 2, 3, 14, 7, 15], pixel-level representations can be subjective, as different doctors may annotate the details of the same object differently [40, 1, 2].
-In OCTA imaging, retinal blood vessels exhibit a uniform color change and have a shape closely resembling a path.
-They can be represented using coordinates along with road width.
-Space Colonization Algorithm (SCA) [50] just meets this requirement.
+In traditional models such as GAN+SLSM [20, 21, 23, 28, 29, 31] and standalone SLSMs [2, 3, 7, 8, 9, 10, 12, 15, 16, 17, 48, 49, 50, 51], pixel-level segmentations are prone to inter-observer variability [1, 2, 44].
+However, in OCTA, retinal vessels have consistent visual traits, uniform grayscale gradients and path-like shapes, making them suitable for coordinate-based modeling with associated widths.
+We adopt the Space Colonization Algorithm (SCA) [54] to represent vessels as structured graphs, enhancing topological coherence and connectivity.
+To bridge the domain gap between initial vessel maps and real OCTA data, we apply a CycleGAN-based refinement module [39], aligning generated structures with realistic image characteristics while preserving topology.
+Unlike prior methods that segment post-rendered images, we introduce a DFS-based segmentation approach [56] applied directly to graph-structured (coordinates, radius) data.
+This intercepts the generation process to ensure anatomical fidelity, governed by a radius threshold $R_{min}$ derived from biostatistical vessel metrics [34, 35, 36, 37].
+3 Methodology
+Figure 1: The architecture of X-GAN where red indicates that X-GAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
+The complete framework of X-GAN is illustrated in Fig. 1 and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
+3.1 Vessel Structure Refinement Module
 Rather than generating pixel-wise vessel segmentations, we formulates the vascular network as a structured graph representation, ensuring topological consistency and robust connectivity.
-The graph consists of the following components, shown in Eq.
-1:
+The graph consists of the following components, shown in Equation 1:
 $V=\left\{(x_{i},y_{i},z_{i}|r_{i})\right\}_{i=1}^{N}$ (1)
@@
 To mitigate this gap, we employ a tuned GAN based on CycleGAN [39] as a structural refinement module, aligning vessel representations with real OCTA distributions while preserving vascular topology.
-We introduce the Depth-First Search (DFS, a nineteenth century algorithm)-based segmentation approach [52], applying DFS directly to structured (coordinates, radius) data before image rendering.
-Unlike GAN+SLSM [17, 26, 21, 27, 28, 19], which segments generated images mixed with real data, our method intercepts the GAN generation process, ensuring segmentation aligns with biostatistical vessel radii.
-The DFS threshold, $R_{min}$, is defined as the minimum radius ratio of main vessels, following established biostatistical principles [29, 30, 31, 33, 32].
-Also, we introduce GSS-RetVein, a dataset capturing rich details of both main vessels and capillaries.
-The dense capillaries of our dataset introduce noise, effectively testing model robustness for SLSMs.
-This detail aids vascular reconstruction, blood flow prediction, and glaucoma progression simulation by enhancing edge and endpoint learning [1, 2, 3].
-Inspired by Harvard glaucoma dataset series [37, 38, 36, 39, 41, 40], we integrate fairness measures to minimize bias and ensure equitable model performance.
-3 Methodology
-The complete framework of GlaGAN is illustrated in Fig.
-1 where red indicates that GlaGAN does not utilize these things in Segmentor $S$, whereas GAN+SLSMs do.
-and it consists of two parts: Vessel Structure Refinement Module Module and Segmentor.
-3.1 Vessel Structure Refinement Module
 Given a vessel structure $X_{s}$ via SCA, our generator $G$ learns to transform it into a realistic OCTA-like representation $X_{r}^{\prime}$, expressed as $G(X_{s})$, while $X_{r}^{\prime}$ undergoes style adaptation to align with the contrast and noise characteristics of real OCTA images $X_{r}$.
 To enhance structural consistency, we modify the original CycleGAN architecture by removing the inverse generator and incorporating a segmentation consistency loss via the segmentor $S$.
-This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Eq.
-2:
+This modification prevents CycleGAN from altering the vessel topology, ensuring that vessel segmentation remains biologically meaningful, as illustrated in Equation 2:
 $L_{\text{seg}}=\mathbb{E}_{X_{r}}\left[\|S(X_{r}^{\prime})-S(X_{r})\|_{1}\right]$ (2)
@@
 $L_{\text{total}}$ $=L_{\text{GAN}}+\lambda_{\text{GP}}L_{\text{GP}}+\lambda_{\text{seg}}L_{\text{seg}}$ (5)
-where Eq.
-3 is the Wasserstein Adversarial Loss, Eq.
-4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Eq.
-5 is the Structural Consistency via Segmentation Loss.
+where Equation 3 is the Wasserstein Adversarial Loss, Equation 4 is the Gradient Penalty for Stability ($\hat{X}$ is the reference vessel representation) and Equation 5 is the Structural Consistency via Segmentation Loss.
 Upon training convergence, the generator produces a vessel structure that not only exhibits realistic contrast properties but also preserves anatomical coherence, ensuring fidelity to real OCTA characteristics.
@@
 First, we construct a graph $G=(V,E)$, where nodes $V$ represent vessel centerline points and edges $E$ connect adjacent vessel points based on local vessel connectivity.
-A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Eq.
-6:
+A vessel segment $e_{ij}$ between nodes $v_{i}$ and $v_{j}$ (the direction is from $i$ to $j$) is assigned a weight, as shown in Equation 6:
 $w(e_{ij})=r_{i}$ (6)
@@
 To extract primary vessels, we apply radius thresholding and DFS traversal.
-For radius filtering, we define a minimum main vessel radius $R_{min}$ and retain only Eq.
-7:
+For radius filtering, we define the ratio of the smallest main vessel radius to the largest main vessel radius $R_{min}$ and retain only Equation 7:
 $V_{\text{main}}=\left\{(x_{i},y_{i},z_{i}|r_{i})\mid r_{i}\geq(R_{\min}*r_{max})\right\}$ (7)
-For DFS traversal, it consists of three steps:
-  • select the optic disc region as the root node
-  • recursively traverse the largest connected component using DFS
-  • stop when no further vessel segments satisfy the radius constraint
+For DFS traversal, it consists of three steps: (1) select the optic disc region as the root node; (2) recursively traverse the largest connected component using DFS; (3) stop when no further vessel segments satisfy the radius constraint.
 When all three steps finished, construct final extracted vessel structure $G_{main}$.
 This approach effectively retains only the primary vascular structure, removing small capillaries and noise.
-This method can be summarized as the following Algorithm 1.
-Algorithm 1 Primary Vessel Extraction via DFS
-0:  Vessel Structure $V=\{(x_{i},y_{i},z_{i}\mid r_{i})\}_{i=1}^{N}$
-0:  Extracted Primary Vessel Graph $G_{\text{main}}$
-1:  Step 1: Vessel Graph Construction
-2:  for all vessel point $(x_{i},y_{i},z_{i}\mid r_{i})\in V$ do
-3:     Identify Neighboring Vessel Points to Form Edge Set $E$
-4:     Assign Edge Weight $w(e_{ij})=r_{i}$
-5:  end for
-6:  Step 2: Vessel Filtering
-7:  Apply Radius Threshold $R_{\text{min}}$
-8:  $V_{\text{main}}=\{(x_{i},y_{i},z_{i}\mid r_{i})\mid r_{i}\geq R_{\text{min}}\cdot r_{\max}\}$
-9:  Remove Weakly Connected Components
-10:  Step 3: DFS-Based Traversal
-11:  Initialize DFS from Optic Disc Region
-12:  for all point $(x_{i},y_{i},z_{i}\mid r_{i})\in V_{\text{main}}$ do
-13:     Recursively Explore Vessel Network via DFS
-14:     Preferentially Expand along Thicker Vessels
-15:  end for
-16:  Step 4: Output
-17:  return  $G_{\text{main}}$
-The key aspect of Algorithm 1 is the threshold $R_{min}$.
-Instead of the final image, we analyze intermediate outputs $S(X_{r})$, CSV data (Fig.
-2(e)), and GAN-derived radius statistics (Fig.
-2(d)).
-For a sample image (Fig.
-2(a)), we computed a Mean (85.31) and Standard Deviation (88.67), determining $R_{min}$ as 0.1968 $r_{max}$.
-Across datasets, values were 0.2011 (GSS-RetVein), 0.2102 (OCTA-500), and 0.1982 (ROSE), standardized to 0.2 for consistency.
-Ophthalmology studies [29, 30, 31, 33, 32] confirm 18-26% capillary coverage in OCTA images, validating this threshold.
-(a) GSS-RetVein
-(b) 3D Visualization
-(c) Statistical Distribution
-(d) Radius Statistics
-(e) Vascular Data
-Figure 2: 3D Visualization of the Pixel Distribution of One Sample, and its Gaussian Statistical Distribution
 4 Dataset and Implementation
 4.1 Dataset
-4.1.1 GSS-RetVein
-GSS-RetVein is a unique and pioneering dataset focused exclusively on retinal vascularity.
-The dataset features an exceptionally clear distribution of capillaries.
-It comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race.
+4.1.1 OCTA-500
+OCTA-500 [1] focuses on retinal blood vessel research, providing $\times6mm$ (300 images, ID range 10001 to 10300) and $\times3mm$ (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies.
+It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
+4.1.2 ROSE
+ROSE [4] is an open-source collection designed for retinal blood vessel segmentation using OCTA images.
+It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a $\times3mm$ foveal-centered area with a resolution of $\times304$ pixels, providing both centerline-level and pixel-level vessel annotations.
+ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a $\times3mm$ area, resized to $\times840$ pixels, with centerline-level annotations.
+(a) UTSW Data
+(b) OCTA-500
+(c) ROSE
+Figure 2: Comparison of UTSW with Other Datasets
+4.1.3 UTSW
+our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki.
+Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
+As show in Fig.
+2, original images come from the latest Intalight device [^1] https://intalight.com/, which is currently the most advanced equipment.
+As shown in Fig.
+2(a), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy.
+UTSW comprises 550 images, both 2D and 3D, from 250 subjects, ensuring fair representation across gender and race.
 We collected retinal vascular data for both eyes (OS and OD) of each subject, with a gender split of 48.8% male and 51.2% female.
@@
 Additional data includes age (47.1 $\pm$ 24.8 years) and examination dates from 2023 to 2024.
-GSS-RetVein’s original images come from the latest Intalight device [^1] https://intalight.com/, offering superior detail compared to OPTOVUE [^2] https://www.visionix.com/ (Fig.
-3).
-As shown in Fig.
-3(a) and Fig.
-3(c), Intalight images feature clearer vessel boundaries and capillary distributions, enhancing segmentation accuracy.
 In supervised learning, their minimal pixel variation introduces valuable noise for testing model robustness [3].
 For 3D reconstruction and evolution prediction, these high-resolution images aid in vascular modeling and blood flow analysis [2], making Intalight data ideal for both tasks.
-(a) Intalight
-(b) OPTOVUE
-(c) Intalight
-(d) OPTOVUE
-Figure 3: Comparison of Data Collected by Different Devices for the Same Patient
-For GSS-RetVein, we eliminate red and green baselines from Intalight device images through a multi-step process.
-First, we apply a dark channel algorithm [5] to remove light artifacts, followed by unsharp masking [6] to enhance high-frequency details.
-The final images (Fig.
-4(a)) undergo manual verification for quality assurance.
-As shown in Fig.
-4, while main vessels remain visible across datasets, magnification reveals that OCTA-500 [1] and ROSE [4] blur capillaries and vessel ends.
-Even the unprocessed baseline image (Fig.
-LABEL:cid) offers superior vessel clarity compared to other datasets (Fig.
-4(c), Fig.
-4(b)).
-Among them, Fig.
-3(a) and Fig.
-3(b) are images of deep angio, and Fig.
-3(c) and Fig.
-3(d) are images of superficial angio.
-They all are images of inner retina vessel.
-(a) GSS-RetVein
-(b) OCTA-500
-(c) ROSE
-Figure 4: Comparison of GSS-RetVein with Other Datasets
-For the representation of 3D imaging of GSS-RetVein, as shown in Fig.
-5, capillaries stand out from the thicker main blood vessels by attaching to them and narrowing at their ends (Fig.
-5(c)).
-The elaborate architecture of blood vessels and the intertwining capillaries contribute to an extremely intricate vascular layout.
-3D imaging part of GSS-RetVein is formatted as a JSON file, which includes 3D X, Y, and Z coordinates and the corresponding radius pixel distribution values.
-For annotation (2D, image format: PNG), our team includes two glaucoma specialists who handle labeling, verification, and manual evaluation, utilizing the Labelme [^3] https://github.com/wkentaro/labelme.
-(a) Top View
-(b) Front View
-(c) Detail
-Figure 5: 3D Imaging of Retinal Vessel from Various Angles
-In general, GSS-RetVein is a dataset that combines 2D and 3D images with extremely high image resolution, and is particularly dedicated to retinal vascular segmentation.
-Table 1: Comparison GlaGAN With Other Models on GSS-RetVein, OCTA-500 and ROSE ($\times$100%)
-Model Dataset
-GSS-RetVein OCTA-500 ROSE
-6mm 3mm 6mm 3mm 3mm
-IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
-U-Net 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
-U-Net++ 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
-Attention U-Net 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
-Mask R-CNN 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
-YOLOv11-x 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
-MedSAM 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
-CauSSL 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
-UniverSeg 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
-S2VNet 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
-Tyche 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
-GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
-Table 2: Conducting Mixed Validation on Different Models using Mixed Datasets ($\times$100%)
-Model Dataset
-ROSE GSS-RetVein OCTA-500 GSS-RetVein (Test) OCTA-500 (Test)
-GSS-RetVein OCTA-500 ROSE OCTA-500 (Train) GSS-RetVein (Train)
-3mm 6mm 3mm 6mm 3mm
-IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
-U-Net 98.34 99.16 98.32 99.15 93.89 96.85 94.23 97.03 96.21 98.07 95.23 97.56 96.98 98.47 96.81 98.38
-U-Net++ 98.99 99.49 98.78 99.39 94.08 96.95 94.46 97.15 95.73 97.82 97.01 98.51 97.59 98.78 96.93 98.44
-Attention U-Net 98.54 99.26 98.23 99.11 94.07 96.94 94.52 97.18 96.99 98.47 96.09 98.01 97.21 98.59 96.78 98.36
-Mask R-CNN 96.37 98.15 96.31 98.12 90.73 95.14 91.89 95.77 94.23 97.03 94.11 96.97 96.68 98.31 95.96 97.94
-YOLOv11-x 98.88 99.44 98.72 99.36 94.11 96.97 94.49 97.17 97.32 98.64 96.54 98.24 96.12 98.02 96.63 98.29
-MedSAM 99.04 99.52 98.98 99.49 95.01 97.41 95.42 97.66 97.54 98.75 95.87 97.89 98.23 99.11 97.46 98.72
-CauSSL 96.83 98.38 95.89 97.90 96.98 98.47 97.83 98.88 99.12 99.56 98.81 99.40 94.87 97.33 95.34 97.62
-UniverSeg 98.96 99.48 98.56 99.28 94.23 97.03 94.59 97.24 96.81 98.38 96.32 98.16 97.91 98.96 98.12 99.05
-S2VNet 99.23 99.61 98.79 99.39 93.89 96.85 95.22 97.55 96.95 98.44 95.83 97.87 97.87 98.95 97.87 98.95
-Tyche 98.78 99.38 98.43 99.21 94.22 97.02 94.98 97.46 96.88 98.41 96.11 98.02 97.84 98.94 97.68 98.87
-GlaGAN 99.55 99.78 99.21 99.60 96.98 98.46 97.69 98.83 98.87 99.43 98.79 99.39 99.58 99.79 99.41 99.70
-Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki.
-Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
-4.1.2 OCTA-500
-OCTA-500 [1] focuses on retinal blood vessel research, providing 6mm×$\times$6mm (300 images, ID range 10001 to 10300) and 3mm×$\times$3mm (200 images, ID range 10301 to 10500) fields of view for wide-field vessel analysis and high-resolution microvascular studies.
-It includes detailed annotations for arteries, veins, large vessels, capillary networks, and 2D/3D Foveal Avascular Zone, facilitating analysis of retinal vessels structures and pathological changes.
-4.1.3 ROSE
-ROSE [4] is an open-source collection designed for retinal blood vessel segmentation using OCTA images.
-It consists of two subsets: ROSE-1 includes 117 OCTA images from 39 subjects, covering a 3mm×$\times$3mm foveal-centered area with a resolution of 304$\times$304 pixels, providing both centerline-level and pixel-level vessel annotations.
-ROSE-2 contains 112 OCTA images from 112 eyes, focusing on the superficial vascular complex within a 3mm×$\times$3mm area, resized to 840$\times$840 pixels, with centerline-level annotations.
-4.2 Implement
-Table 3: Comparison Vessel Structure Refinement Images with Real OCTA Images ($\times$100%)
-Model Dataset
-GSS-RetVein OCTA-500 ROSE
-6mm 3mm 6mm 3mm 3mm
-SSIM MSE SSIM MSE SSIM MSE SSIM MSE SSIM MSE
-GAN 91.76 5.69 90.39 4.43 89.69 9.85 91.06 7.36 87.84 8.73
-CycleGAN 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
-GAN 93.62 3.29 93.87 3.79 91.28 6.98 93.19 5.54 91.73 4.18
-CycleGAN 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
+Our team includes two attending glaucoma physicians who are responsible for the detailed labeling of the main retinal blood vessels using the tool Labelme [^2] https://github.com/wkentaro/labelme.
+4.2 Implementation
 4.2.1 Hyperparameter
-for GlaGAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70).
+For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN ($\times70$).
 The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$.
-GlaGAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation.
-Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
-4.2.2 Hardware
-the hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
-4.2.3 Evaluation Metric
+X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation.
+Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$) and flipping.
+4.2.2 Evaluation Matrix
 We use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth.
@@
 While SSIM aligns with human perception, MSE focus on absolute pixel differences.
-Table 4: Ablation Experiment of GlaGAN ($\times$100%)
+4.2.3 Experimental Setup
+We selected the following deep learning models as reference benchmarks: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [15], UniverSeg [16], S2VNet [17] and Tyche [19].
+These models are leaders in the fields of image segmentation and medical image processing.
+For their parameter adjustment, we use Optuna [^3] https://optuna.org/ to optimize their learning green, decay and other hyperparameters.
+For the training and testing of the model, we use ten cross-validation methods.
+The evaluation indicators are finally averaged.
+4.2.4 Hardware
+The hardware for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
+5 Experimental Result
+5.1 Comparative Experiment
+Table 1: Comparison X-GAN with Other Models on Different Datasets ($\times$100%)
 Model Dataset
-GSS-RetVein OCTA-500 ROSE
+UTSW OCTA-500 ROSE
 6mm 3mm 6mm 3mm 3mm
 IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
-GAN U-Net 92.89 96.31 92.44 96.07 93.31 96.56 92.84 96.31 93.37 96.59
-GAN U-Net++ 93.62 96.70 92.87 96.30 93.72 96.73 93.03 96.36 92.84 96.31
-GAN Attention U-Net 93.56 96.67 93.12 96.44 93.65 96.70 92.85 96.31 93.42 96.61
-GAN MedSAM 93.73 96.76 92.76 96.24 92.83 96.30 92.47 96.10 93.76 96.75
-GAN DFS 95.81 97.86 95.24 97.56 94.83 97.33 94.87 97.35 94.73 97.27
-CycleGAN U-Net 93.44 96.61 93.42 96.60 94.73 97.26 93.97 96.89 93.76 96.75
-CycleGAN U-Net++ 93.72 96.76 93.62 96.70 94.58 97.18 94.36 97.05 93.19 96.46
-CycleGAN Attention U-Net 93.79 96.80 93.45 96.61 94.59 97.18 94.29 97.01 94.03 96.89
-CycleGAN MedSAM 94.02 96.92 94.37 97.10 93.83 96.72 94.37 97.19 94.18 96.96
-CycleGAN DFS 96.23 98.08 96.42 98.18 96.68 98.33 96.27 98.13 95.83 97.84
-CycleGAN U-Net 94.56 97.20 94.52 97.18 94.86 97.31 94.79 97.27 94.19 96.94
-CycleGAN U-Net++ 94.98 97.43 94.36 97.10 94.83 97.29 95.09 97.41 94.27 96.98
-CycleGAN Attention U-Net 95.06 97.47 95.12 97.50 95.07 97.39 94.97 97.33 94.39 97.04
-CycleGAN MedSAM 95.33 97.61 95.21 97.55 95.36 97.55 95.11 97.41 94.82 97.27
-GlaGAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
-5 Experimental Result
-5.1 Experimental Setup
-We selected many SOTA models for experimental verification, which are: U-Net [10], U-Net++ [7], Attention U-Net [12], Mask R-CNN [9], YOLOv11-x [11], MedSAM [8], CauSSL [13], UniverSeg [14], S2VNet [15], Tyche [16], GAN [34] and CycleGAN [35].
-Among them, we used tool Optuna [^4] https://optuna.org/ to optimize their parameters.
-All final experimental results are the average of ten cross-validation times.
-* means models’ hyperparameters have been optimized by us.
-5.2 Comparative Experiment
-As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, GlaGAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation.
-As shown in Table 2, the second row is the test data set, and the third row is the training data set.
-Models trained on GSS-RetVein and OCTA-500 were tested on ROSE, where GlaGAN achieved the highest IoU and Dice scores, outperforming S2VNet, CauSSL, and MedSAM.
-For generalization, training on ROSE and testing on GSS-RetVein and OCTA-500 confirmed GlaGAN’s superiority, followed by MedSAM and CauSSL.
-Across different scanning resolutions, GlaGAN consistently led in segmentation accuracy, with MedSAM and UniverSeg performing well, while U-Net variants lagged.
-These results affirm GSS-RetVein’s quality and GlaGAN’s segmentation excellence.
-5.3 Ablation Experiment
-For baseline GAN model, GAN and CycleGAN are selected.
-For baseline segmentor, U-Net, U-Net++, Attention U-Net and MedSAM are selected.
-For GlaGAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy.
-As shown in Table 4, since main blood vessels make up a small proportion, the value calculated is also relatively low.
-When the minimum radius ratio is found, like GlaGAN, aligning with biostatistical properties [29, 30, 31, 33, 32], segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
-Table 5: Ablation Experiment of Hyperparameters ($\times$100%)
-Hyperparameter Dataset
-GSS-RetVein OCTA-500 ROSE
-6mm 3mm 6mm 3mm 3mm
-IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
-Rmin 0.00 train 1.03 2.04 1.00 1.98 1.31 2.59 1.22 2.41 1.13 2.23
-test 0.75 1.49 0.87 1.73 0.91 1.80 0.77 1.52 0.93 1.84
-0.10 train 41.32 58.47 41.98 59.12 43.09 60.23 45.57 62.59 44.29 61.38
-test 40.29 57.42 42.75 59.89 42.31 59.45 45.06 62.13 43.97 61.08
-0.20 train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
-test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
-0.30 train 87.15 93.13 86.29 92.64 86.04 92.49 85.74 92.30 84.37 91.52
-test 84.17 91.40 83.96 91.27 84.09 91.35 84.02 91.31 83.98 91.28
-0.40 train 57.92 73.34 56.44 72.15 55.86 71.65 55.45 71.28 57.09 72.69
-test 58.73 73.98 57.85 73.29 57.29 72.85 56.93 72.53 57.24 72.80
-0.50 train 55.72 71.56 56.98 72.60 54.79 70.78 56.93 72.53 57.04 72.65
-test 54.17 70.25 53.83 69.99 52.62 68.97 53.71 69.89 53.83 69.99
-0.60 train 40.26 57.40 41.27 58.42 40.39 57.53 41.03 58.17 39.27 56.39
-test 35.62 52.51 38.92 56.03 37.46 54.50 37.08 54.10 35.74 52.63
-0.70 train 35.36 52.25 34.72 51.55 33.91 50.64 37.06 54.08 35.94 52.85
-test 32.24 48.76 31.83 48.27 32.64 49.21 33.39 50.06 34.07 50.83
-0.80 train 17.83 30.26 19.25 32.29 18.83 31.69 17.79 30.21 16.68 28.59
-test 26.37 41.73 25.73 40.93 25.49 40.61 25.13 40.15 24.80 39.73
-0.90 train 9.85 17.94 10.32 18.71 9.89 17.99 11.27 20.26 9.91 18.03
-test 12.24 21.82 13.06 23.10 12.94 22.92 11.98 21.39 10.57 19.12
-1.00 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-Node 0 train 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-test 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-5000 train 6.34 11.92 7.81 14.49 6.28 11.82 6.43 12.09 7.02 13.12
-test 6.06 11.43 5.86 11.07 5.93 11.19 6.18 11.64 6.58 12.35
-10000 train 17.82 30.24 18.92 31.81 17.96 30.44 19.06 32.02 17.48 29.75
-test 20.59 34.15 21.76 35.73 22.84 37.19 21.96 36.00 20.98 34.69
-15000 train 20.58 34.14 21.96 36.00 20.84 34.50 21.37 35.21 20.71 34.29
-test 28.73 44.63 29.64 45.71 30.74 47.04 28.93 44.87 29.74 45.83
-20000 train 41.47 58.61 42.34 59.49 41.65 58.80 42.07 59.18 44.03 61.14
-test 42.06 59.17 41.11 58.26 41.93 59.04 40.62 57.77 41.98 59.12
-25000 train 49.38 66.12 52.14 68.53 50.39 66.96 52.97 69.26 51.07 67.60
-test 52.01 68.41 51.49 67.98 51.76 68.21 52.44 68.79 53.06 69.35
-30000 train 59.23 74.39 59.17 74.34 58.48 73.78 57.99 73.40 58.49 73.79
-test 61.25 75.97 60.84 75.64 59.73 74.78 61.79 76.37 60.83 75.63
-35000 train 62.33 76.79 63.34 77.56 62.15 76.64 61.96 76.49 62.15 76.64
-test 62.51 76.96 64.08 78.12 63.24 77.47 62.44 76.89 63.19 77.43
-40000 train 62.74 77.13 64.27 78.25 63.74 77.88 63.96 78.06 63.49 77.64
-test 62.59 77.03 64.13 78.16 63.96 78.06 61.84 76.39 63.97 78.07
-45000 train 63.37 77.57 65.03 78.81 65.17 79.57 65.03 78.81 65.93 79.41
-test 65.72 79.30 66.34 79.77 64.44 78.39 65.39 78.95 66.67 80.00
-50000 train 66.53 79.89 66.02 79.53 65.52 79.14 65.74 79.31 66.03 79.54
-test 69.34 81.89 68.18 81.06 69.01 81.65 66.95 80.19 68.39 81.24
-50000 train 71.24 83.20 72.47 84.02 69.08 81.71 68.79 81.52 69.28 81.83
-test 73.69 84.85 72.61 84.12 74.17 85.15 75.36 85.94 74.13 85.10
-55000 train 75.48 85.99 76.24 86.54 76.14 86.47 75.83 86.25 74.19 85.17
-test 76.02 86.38 75.47 85.98 74.45 85.34 73.99 85.05 74.73 85.57
-60000 train 79.28 88.47 78.96 88.23 79.48 88.61 77.20 87.13 75.96 86.35
-test 78.49 87.94 77.85 87.52 77.02 87.03 79.36 88.52 75.89 86.29
-65000 train 85.27 92.05 86.09 92.52 85.78 92.29 85.96 92.42 86.06 92.50
-test 84.18 91.40 84.07 91.33 83.33 90.89 82.39 90.37 83.09 90.76
-70000 train 97.46 98.71 97.28 98.62 97.08 98.52 96.96 98.46 97.07 98.51
-test 98.32 99.15 98.29 99.14 99.03 99.51 99.01 99.50 98.76 99.37
-…… train 98.71 99.35 98.07 99.03 98.86 99.43 98.14 99.06 98.05 99.02
-test 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
-As shown in Table 5, the optimal choice of $R_{min}$ (0.2) enables GlaGAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%.
-When $R_{min}=1$, GlaGAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold.
-When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
-Also, as shown in Fig.
-ref1 and Table 5, a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy.
-Given that the R values of GlaGAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
-Consequently, their curves largely overlap, with GlaGAN exhibiting a slight increase.
-5.4 Evaluation of Vessel Structure Refinement
+U-Net [10] 95.14 97.51 95.23 97.56 96.93 98.44 98.75 99.37 97.94 98.96
+U-Net++ [7] 95.68 97.81 96.64 98.29 97.65 98.81 97.96 98.97 97.52 98.75
+Attention U-Net [12] 95.78 97.84 96.04 97.98 97.05 98.50 98.61 99.30 97.92 98.95
+Mask R-CNN [9] 92.35 96.02 91.35 95.48 92.06 95.87 93.50 96.64 91.23 95.41
+YOLOv11-x [11] 95.81 97.86 96.55 98.24 95.93 97.92 97.66 98.82 97.83 98.89
+MedSAM [8] 96.38 98.16 96.47 98.20 96.01 97.95 98.94 99.47 97.89 98.93
+CauSSL [15] 95.23 97.56 95.47 97.68 96.32 98.04 98.15 99.08 97.80 98.89
+UniverSeg [16] 95.30 97.58 95.52 97.70 96.58 98.17 98.42 99.22 97.07 98.01
+S2VNet [17] 95.68 97.81 95.36 97.63 96.45 98.11 98.27 99.15 97.92 98.95
+Tyche [19] 95.80 97.85 95.19 97.53 96.28 98.02 98.11 99.06 97.76 98.87
+X-GAN 99.41 99.71 98.66 99.33 99.21 99.60 99.42 99.71 99.19 99.59
+As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, X-GAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation.
 As shown in Table 2, comparison between CycleGAN [39] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%.
@@
 The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
+Table 2: Comparison Vessel Structure Refinement with Real OCTA ($\times$100%)
+Model Dataset
+UTSW OCTA-500 ROSE
+6mm 3mm 6mm 3mm 3mm
+SSIM MSE SSIM MSE SSIM MSE SSIM MSE SSIM MSE
+CycleGAN [39] 93.75 2.36 94.32 2.39 93.56 1.62 94.58 4.46 91.01 3.83
+CycleGAN (ours) 95.21 1.88 95.89 1.85 94.82 1.73 96.04 3.10 95.32 2.16
+Human Evaluation: the medical team verified the corrected images.
+Luckily, since this study only aimed to segment only the main blood vessels, the small size of the capillaries caused errors in radius generation, which affected the accuracy of the indicators.
+5.2 Ablation Eperiment
+For baseline GAN model, GAN [38] and CycleGAN [39] are selected.
+For baseline segmentor, U-Net [10], U-Net++ [7], Attention U-Net [12] and MedSAM [8] are selected.
+For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy.
+As shown in Table 3, since main blood vessels make up a small proportion, the value calculated is also relatively low, aligning with biostatistical properties [34, 35, 36, 37], segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
+Table 3: Ablation Experiment of X-GAN ($\times$100%)
+Model Dataset
+UTSW
+6mm 3mm
+IoU Dice IoU Dice
+GAN [38] + U-Net [10] 92.89 96.31 92.44 96.07
+GAN [38] + U-Net++ [7] 93.62 96.70 92.87 96.30
+GAN [38] + Attention U-Net [12] 93.56 96.67 93.12 96.44
+GAN [38] + MedSAM [8] 93.73 96.76 92.76 96.24
+GAN [38] + DFS ($R_{min}$=0.2) 95.81 97.86 95.24 97.56
+CycleGAN [39] + U-Net [10] 93.44 96.61 93.42 96.60
+CycleGAN [39] + U-Net++ [7] 93.72 96.76 93.62 96.70
+CycleGAN [39] + Attention U-Net [12] 93.79 96.80 93.45 96.61
+CycleGAN [39] + MedSAM [8] 94.02 96.92 94.37 97.10
+CycleGAN [39] + DFS ($R_{min}$=0.2) 96.23 98.08 96.42 98.18
+CycleGAN (ours) + U-Net [10] 94.56 97.20 94.52 97.18
+CycleGAN (ours) + U-Net++ [7] 94.98 97.43 94.36 97.10
+CycleGAN (ours) + Attention U-Net [12] 95.06 97.47 95.12 97.50
+CycleGAN (ours) + MedSAM [8] 95.33 97.61 95.21 97.55
+X-GAN 99.41 99.71 98.66 99.33
+As shown in Fig.
+3(a), the optimal choice of $R_{min}$ (0.2) enables X-GAN to significantly outperform other models in segmentation, which is so high that it is nearly perfect, approaching 100%.
+When $R_{min}=1$, X-GAN segments only the largest vessels, resulting in an extremely low segmentation index as only the starting points meet the threshold.
+When $R_{min}$ is 0, it is equivalent to no filtering, directly the original image.
+(a) $R_{min}$
+(b) The Number of Nodes NN
+Figure 3: Ablation Experiments of Key Parameters ($\times$100%)
+Also, as shown in Fig.
+3(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy.
+Given that the R values of X-GAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
 5.3 Efficiency Evaluation
-As shown in Table 6, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity.
+Compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity.
 It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
-Table 6: Comparison of Performance Indicators between DFS and SLSMs ($\times$100%)
-Indicator SLSMs DFS (ours)
-GPU $\checkmark$ $\times$
-Annotation $\checkmark$ $\times$
-Training $\checkmark$ $\times$
-Explainability $\times$ $\checkmark$
-2D and 3D $\times$ $\checkmark$
-Generalization $\times$ -
+Table 4: Comparison of Performance Details between DFS (our) and SLSMs
 Model Details
 Inference Time Parameters FLOPs GPU Utilization
-U-Net 21ms 31M 98 low
-U-Net++ 62ms 76M 304 high
-Attention U-Net 83ms 62M 215 slightly high
-MedSAM 425ms 109M 1250 extremely high
-DFS (CPU) 0.018ms 0M 0.00028 -
-DFS (GPU) $\leq$ 0.01ms 0M 0.00028 extremely low
-Table 7: Comparison of Performance Details between DFS and SLSMs
+U-Net [10] 21ms 31M 98 low
+U-Net++ [7] 62ms 76M 304 high
+Attention U-Net [12] 83ms 62M 215 slightly high
+MedSAM [8] 425ms 109M 1250 extremely high
+DFS (ours) 0.018ms 0M 0.00028 -
+DFS (ours, GPU based) $\leq$ 0.01ms 0M 0.00028 extremely low
 Compared to the segmentor, our adjusted DFS offers unmatched advantages, as shown in Table 4.
@@
 As shown in Fig.
-6, 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis.
+4, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis.
 In Fig.
-6(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow.
+4(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow.
 DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries.
 Extending this to 3D, we achieve precise main vessel extraction, with Fig.
-6(e) preserving vessel integrity and offering a clearer structure than the original image (Fig.
-6(c)).
+4(c) preserving vessel integrity and offering a clearer structure than the original image (Fig.
+4(b)).
 (a) Segmentation of Main Retinal Vessels (2D)
-(b) Original Image (2D)
-(c) Original Image (3D)
-(d) Segmentation (2D)
-(e) Segmentation (3D)
-Figure 6: Segmentation Results of GlaGAN on GSS-RetVein (2D and 3D)
+(b) Original Image (3D)
+(c) Segmentation (3D)
+Figure 4: Segmentation of X-GAN (2D and 3D)
 5.5 Analysis
-Optimal selection of $R_{min}$ values allows GlaGAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1, Table 2, Table 4).
-Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [29, 30, 31, 33, 32].
-We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted GlaGAN to achieve amazing segmentation results.
-Unlike GAN+SLSMs or SLSMs, GlaGAN does not require additional data to avoid fitting issues.
+Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1 and Table 3).
+Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [34, 35, 36, 37].
+We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results.
+Unlike GAN+SLSMs or SLSMs, X-GAN does not require additional data to avoid fitting issues.
 Its segmentor is efficient, computationally friendly, and free from pixel mapping errors associated with SLSMs segmentation.
-To the best of our knowledge, GlaGAN is the best model currently, tailored to segment OCTA retinal main vessel.
+Studies [32, 33] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models.
 6 Conclusion
-In this paper, we propose GlaGAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs.
-We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation.
-Extensive experiments demonstrate that GlaGAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability.
+In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs.
+Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability.
 Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
-Acknowledgment
-This research is partially supported by NASA under grant 80NSSC22K0144, UTSW GMO No. 241213, and NIH grant 1R01AG083179-01.
+7 Future Work
+With access to state-of-the-art glaucoma detection equipment, we have curated a high-quality image dataset featuring precisely annotated retinal vascular images.
+This dataset is well-suited for glaucoma pathology research and computer vision.